### v1

# W-RAG: Weakly Supervised Dense Retrieval in RAG for Open-domain Question Answering

Abstract.
In knowledge-intensive tasks such as open-domain question answering (OpenQA), Large Language Models (LLMs) often struggle to generate factual answers relying solely on their internal (parametric) knowledge. To address this limitation, Retrieval-Augmented Generation (RAG) systems enhance LLMs by retrieving relevant information from external sources, thereby positioning the retriever as a pivotal component. Although dense retrieval demonstrates state-of-the-art performance, its training poses challenges due to the scarcity of ground-truth evidence, largely attributed to the high costs of human annotation. In this paper, we propose W-RAG by utilizing the ranking capabilities of LLMs to create weakly labeled data for training dense retrievers. Specifically, we rerank the top-$K$ passages retrieved via BM25 by assessing the probability that LLMs will generate the correct answer based on the question and each passage. The highest-ranking passages are then used as positive training examples for dense retrieval. Our comprehensive experiments across four publicly available OpenQA datasets demonstrate that our approach enhances both retrieval and OpenQA performance compared to baseline models. Source code is published  [^1] https://github.com/jmnian/weak_label_for_rag.
Retrieval Augmented Generation, Open-domain Question Answering, Weak Supervised Learning, Dense Retrieval
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Open-domain question answering (OpenQA) dating back to the 1960s (Jr. et al., 1961) provides natural language-like answers to reply users’ questions. OpenQA adopts the “Retriever-Reader” architecture (Zhu et al., 2021), where the retriever retrieves relevant passages for the reader to generate answers. Previous studies (Lewis et al., 2020; Xiong et al., 2020a) adopt a seq2seq model as a reader and train it on the labeled dataset. Recently, large language models (LLMs) like GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023) have demonstrated astonishing performance on various tasks, including OpenQA, attributing to the substantial amount of knowledge stored in their internal parameters. Despite the unprecedented achievements of LLMs, they face constraints such as the inability to consistently integrate up-to-date knowledge, as their parametric knowledge is fixed after being trained on huge datasets. Additionally, they are prone to generating plausible but non-factual responses, known as hallucinations (Welleck et al., 2020).
To overcome the limitations of LLMs’ parametric knowledge, retrieval augmented generation (RAG) (Lewis et al., 2020; Guu et al., 2020) is explored, equipping LLMs with a retriever to gather necessary evidence from external sources. Among the two components of RAG, improving the retriever is more feasible due to the recent trend of black-box APIs (OpenAI, 2023) and the high cost and time requirements of fine-tuning open-source LLMs (Dubey et al., 2024). The retriever, a critical part of RAG, is typically either a traditional unsupervised retriever like BM25 (Robertson and Zaragoza, 2009) or a more advanced neural retriever, such as dense retrieval (Nogueira et al., 2020; Xiong et al., 2020b; Khattab and Zaharia, 2020a; Karpukhin et al., 2020b), which encodes questions and passages into the same embedding space and then measures the question-passage relevance score by vector similarity. A key challenge in training dense retrievers is the scarcity of human-annotated data, as in OpenQA, human-labeled evidence passages are often unavailable. Methods like UPR (Sachan et al., 2022a), which ranks passages based on the likelihood of LLMs generating the question given the passage, and AAR (Yu et al., 2023), which combines the top-$K$ passages ranked by the LLM’s averaged cross-attention scores with ground-truth passages as positive passages, have been employed to train dense retrievers. These approaches can train retrievers to find semantically relevant passages but do not guarantee improved RAG performance in OpenQA. As Cuconasu et al. (2024) demonstrated, retrieving relevant passages that cannot answer the question may negatively impact RAG performance in OpenQA tasks. Neither cross-attention scores nor the likelihood of LLMs generating the question based on the input passage explicitly ensures that the question can be answered by the “positive” passage.
To address the scarcity of training data for dense retrievers in RAG for OpenQA, we propose extracting weak labels from existing OpenQA question-answer pairs by leveraging the ranking capabilities of LLMs (Sun et al., 2023). Specifically, we first use BM25 to retrieve the top-$K$ passages for a question, then pair each passage with the question. We rank the passages based on the likelihood that the LLM would generate the question’s ground-truth answer from each question-passage pair. Only the top-ranked passage is selected as the positive example for the question, and we train the dense retrievers using in-batch negative sampling. Our method evaluates the relevance of the question-passage pair by the likelihood of the answer given a passage and the question, making it particularly suitable for OpenQA, where retrieved passages must be capable of providing the correct answer.
We conducted comprehensive experiments across four publicly available OpenQA datasets, and the results demonstrate that our approach enhances both retrieval and OpenQA performances compared to baseline models. Our contributions can be summarized as:
  • We propose W-RAG, a general framework to generate weak labels from question-answer pairs for training dense retrievers in RAG.
  • Comprehensive experiments were conducted, and the results prove that W-RAG improves both retrieval and OpenQA performance over baselines.
  • Open source LLMs were used, and the code was released anonymously to ensure reproducibility.
2. Related Work
2.1. Dense Retrieval
Traditional information retrieval (IR) method are based on exact term matching, like BM25. While it is still widely used due to its efficiency, effectiveness, and robustness, it suffers from the well-known issue of lexical gap (Berger et al., 2000). To address this, leveraging the neural networks, dense retrieval (DR) employs pre-trained language models like BERT (Devlin et al., 2019), to encode questions and passages into embeddings, and measures the question-passage relevance score by the vector similarly in the embedding space. Specifically, DP encodes the whole corpus into embeddings and builds the index, such as Faiss (Douze et al., 2024), on them. When a new question comes in, DR encodes the question into an embedding and performs a nearest neighbor search. DR can be classified into two categories: supervised, like DPR (Karpukhin et al., 2020b), TAS-B (Hofstätter et al., 2021), and ColBERT (Khattab and Zaharia, 2020a; Santhanam et al., 2022); unsupervised, like Contriever(Izacard et al., 2022) and ReContriever(Lei et al., 2023a), based on the training method.
DPR utilizes a dual-tower architecture, with one BERT model dedicated to encode questions and another to encode passages. The similarity between the question and passage embeddings is then calculated, aiming to maximize the log-likelihood of the positive passage. ColBERT uses the same BERT model for both the question and passage encoders, differentiating them by appending a unique special token after the [CLS] token. Unlike DPR, which directly compares question and passage embeddings, ColBERT introduces a late interaction mechanism. It computes the similarity between each question token and all passage tokens, followed by maximum pooling over these similarities. The final similarity score for a question-passage pair is the sum of the pooled scores. TAS-B groups queries based on their embedding similarities and applies a training data sampling technique along with dual-teacher supervision distillation. Contriever trains a bi-encoder model using contrastive learning, generating positive question-passage pairs from an unlabeled corpus. ReContriever follows the same approach as Contriever for generating weak question-passage pairs but adds a self-scoring mechanism during training, where the loss is weighted by these scores.
2.2. RAG for OpenQA
RAG models have been applied to OpenQA, demonstrating significant performance improvement. Different RAG models have been proposed to solve the critical issues in OpenQA, such as how to retrieve relevant passages (Yu et al., 2023; Shi et al., 2023; Shao et al., 2023; Trivedi et al., 2023), when to call the retriever (Jiang et al., 2023b; Wang et al., 2023; Cheng et al., 2024), and how to decrease the computational complexity (Li et al., 2023; Jiang et al., 2023a; Xu et al., 2024; Kim et al., 2024).
Since the retriever is a critical component in RAG, some studies have tried to improve the quality of retrieved passages, including training a better retriever (Yu et al., 2023; Shi et al., 2023) and prompt engineering (Shao et al., 2023; Trivedi et al., 2023). REPLUG (Shi et al., 2023) trains a dense retriever by minimizing the KL divergence between the retrieval likelihood and the language model likelihood computed as the normalized language model probability of the ground-truth, given the question and passage. AAR (Yu et al., 2023) combines the top-$K$ passages ranked by the LLM’s averaged corss-attention scores with ground-truth passages as positive passages and then follow ANCE (Xiong et al., 2021) to sample negative passages to train the dense retriever. ITER-RETGEN (Shao et al., 2023) leverages the model output from the previous iteration as a specific context to help retrieve more relevant knowledge. IRCoT (Trivedi et al., 2023) also adopt a similar approach to perform retrieval, but applies CoT for generating responses.
In some cases, LLMs can generate factual content well without external knowledge, and retrieving passages will decrease RAG’s performance. Self-RAG (Wang et al., 2023) and FLARE (Jiang et al., 2023b) both fine-tune the LLMs to call the search engine automatically when external knowledge is needed. UAR (Cheng et al., 2024) trains classifiers to identify the need for external knowledge.
Applying all the top-$K$ retrieved passages as context not only increases the computational complexity and inference latency but also brings noise. Numerous methods have been proposed to compress the retrieved passages. Selective-Context (Li et al., 2023) filters non-essential lexical units by the summarization of self-information of each token contained in the unit. LongLLMLingua (Jiang et al., 2023a) contrasts the perplexity score of each token in the passage with the perplexity score of the same token conditioned on the question and adopts this conservative perplexity score to filter out tokens. RECOMP (Xu et al., 2024) leverages the summarization models to summarize the retrieved passages. SuRe (Kim et al., 2024) summarizes retrieved passages conditioned on each answer candidates generated by prompting the LLMs and then selects the top summarizations ranked by LLMs through a combination of pointwise and pairwise scoring method.
Figure 1. W-RAG fits into the general RAG pipeline by training the retriever with LLM generated weak labels. Following the steps from top to bottom, we retrieve candidate passages using BM25, present each passage to the LLM to rerank based on the answer likelihood, then use the reranked top passage to train the retriever to finally enhance LLM’s response quality through the standard RAG pipeline. The entire prompt is shown in Figure 2.
2.3. LLMs for Ranking
Concerning the generation of weakly labeled ranking data using LLMs as passage rankers, existing work can be categorized into three categories. The first involves a zero-shot listwise ranking strategy, where a prompt containing the question, instructions, and multiple passages are given to the LLM, and the LLM generates a permutation of ranked indices (Baldelli et al., 2024; Ma et al., 2023; Pradeep et al., 2023; Shen et al., 2023). The second category prompts the LLM with passages and directly instructs it to rate their relevance (Zhuang et al., 2024), or perform pairwise comparisons (Qin et al., 2023). The third approach calculates the question generation likelihood when a passage is included in the prompt (Cho et al., 2023; Sachan et al., 2022b; Zhuang et al., 2023). Our method for scoring passages aligns with this third approach; however, as mentioned in Section 1, we rank the passages by the likelihood of the LLM generating the question’s ground-truth answer based on the question-passage pair.
3. Method
As illustrated in Figure 1, W-RAG introduces a novel approach for training the retrieval component within the RAG pipeline from scratch. Our method leverages a weakly-supervised training paradigm that requires only a set of question-answer pairs and an evidence corpus. The W-RAG process unfolds in three stages: first, it retrieves relevant passages from the evidence corpus; second, it employs a LLM to generate weak labels from them by reranking the retrieved passages by the likelihood of generating the ground-truth answer; finally, the dense retiever is trained using these weak labels. The resulting dense retriever is then capable of retrieving evidence, thereby enhancing the performance of LLMs across a wide array of tasks.
Figure 2. Prompts used for weak label generation and question answering
3.1. Weak-label Generation
Given an evidence corpus $\mathcal{C}=\{d_{1},\dots,d_{N}\}$, we use a simple retreiver to conduct first-stage retrieval. Given a question $q$, the retriever ranks each passage by relevance and selects the top-$K$ passages $\mathcal{P}=\{s_{1},\dots,s_{K}\}$, where $\mathcal{P}\subset\mathcal{C}$. The corpus may comprise chuncked Wikipedia articles or from online sources such as Reddit and Common Crawl, while the question-answer pairs could be sourced from OpenQA platforms such as Quora and StackOverflow. The simple first-stage retriever could be lexical such as BM25 (Robertson and Zaragoza, 2009), or a dense retriever. The primary objective at this stage is to maximize recall within a manageable set of retrieved passages. Ideally, this step curates evidence passages containing the correct answer to the question
W-RAG is motivated by the hypothesis that an autoregressive LLM is more likely to generate the correct answer if the provided passage contains the necessary information to answer the question. Our approach aims to harness signals from the LLM’s downstream tasks to evaluate the effectiveness of a passage in eliciting the correct answer. It is important to note that traditional relevance measures, such as term overlap or semantic similarity, do no necessarily guarantee that the most relevant passage contains the needed information. Instead, our target is on identifying passages that most effectively prompt the LLM to produce the correct answer and using these passages to train the dense retriever. In this context, we redefine passage relevance as the degree to which a passage can elicit the ground-truth answer, and this criterion serves as the target for training the dense retriever.
As depicted in Figure 1 Step 2, we construct weak label generation prompts with candidate passage $s_{i}$, question $q$, some instructions $I$, and ground-truth answer $a$. The downstream signal is captured through the conditional probability of the ground-truth answer, denoted as $p(a|s_{i},q,I)$:
(1) $p(a|s_{i},q,I)=\prod_{j}\text{LLM}(a_{j}|s_{i},q,I,a_{<j})$
For each token in the input prompt, the LLM assigns logits, which after applying a softmax operation, correspond to the probability of each vocabulary token being the next in the sequence. The probability of the $j$th answer token, denoted as $\text{LLM}(a_{j}|s_{i},q,I,a_{<j})$, is directly extracted from the logits of the preceding token. Since the ground-truth answer typically consists of multiple tokens, the cumulative probability can diminish rapidly. To address this, we simply take the logarithm on both sides of Equation 1, converting each token’s probability into log-likelihood. The overall probability of the answer $a$ is then computed as the average of the sum of the log-likelihoods for each token in the answer:
(2) $\log p(a|s_{i},q,I)=\frac{1}{|a|}\sum_{j}\log\text{LLM}(a_{j}|s_{i},q,I,a_{<j})$
We consider this log-likelihood as the relevance score for the candidate passage $s_{i}$. Each of the $K$ candidate passages is weakly labeled in this manner, resulting in a ranked list of passages, denoted as $\mathcal{S}=\{s_{r_{1}},\dots,s_{r_{K}}\}$. This list is produced by sorting the set of candidate passages $\mathcal{S}$ according to their relevance scores.
3.2. Training Dense Retriever
Once we have accumulated a sufficient number of weakly labeled passages for a total of $M$ question-answer pairs, denoted as $\mathcal{W}=\{\mathcal{S}_{1},\dots,\mathcal{S}_{M}\}$, we are ready to train a dense retriever that will assign higher scores to passages more likely to elicit the correct answer. In principle, any dense retriever, regardless of the specific objective function, can be trained using the generated weak labels. In this paper, we investigate two representative dense retrievers: DPR (Karpukhin et al., 2020a) and ColBERT (Khattab and Zaharia, 2020b).
3.2.1. DPR
DPR utilizes a bi-encoder architecture, where the question and passage are independently mapped to an embedding space through two separate BERT encoders. For a given question $q$, the relevance score $R_{q,s}$ for each passage $s$ is given by the similarity between their respective [CLS] token embeddings, $\mathbf{e}_{q}$ and $\mathbf{e}_{s}$. We use cosine similarity as the similarity function:
(3) $R_{q,s}=\cos(\mathbf{e}_{q},\mathbf{e}_{s})$
Following the in-batch negative training method introduced in DPR, we process each ranked list $\mathcal{S}\in\mathcal{W}$ by extracting the top-ranked passage $s_{i}$ along with its associated question $q_{i}$. These pairs are then grouped into batches of size $n$ to form training batches, optimizing the retriever using a Multiple Negatives Ranking (MNR) loss (Henderson et al., 2017). In this approach, for the $i$th question-passage pair, $s_{i}$ is treated as the positive example, while the other passages within the same batch are treated as negative examples. The loss function is thus:
(4) $\mathcal{L}_{MNR}=-\sum_{i=1}^{n}\log(\frac{\exp(\alpha\cdot R_{q_{i},s_{i}})}{\sum_{j=1}^{n}\exp(\alpha\cdot R_{q_{i},s_{j}})})$
The scaler $\alpha$ is used to amplify the cosine similarity score, usually set at 20 according to the default setting in sentence-transformers  [^2] https://sbert.net/docs/package_reference/sentence_transformer/losses.html#multiplenegativesrankingloss.
3.2.2. ColBERT
ColBERT employes a bi-encoder architecture with late interaction, where the question and passage are independently encoded using a shared BERT model to obtain embeddings $E_{q}$ and $E_{s}$. The relevance score for each passage is computed on the question side using a “MaxSim” operation, which sums the maximum similarities between any passage token embedding with each question token embedding:
(5) $R_{q,s}=\sum_{i\in|E_{q}|}\max_{j\in|E_{s}|}\text{dot}(E_{q_{i}},E_{s_{j}})$
Here, $\text{dot}(x,y)$ represents the dot product between $x$ and $y$. The training dataset for ColBERT consists of triplet samples $\langle q,s^{+},s^{-}\rangle$, extracted from each ranked list $\mathcal{S}$. In these triplets, $s_{r_{1}}$ is treated as the positive passage $s^{+}$, while $s_{r_{2}},s_{r_{3}},\dots,s_{r_{m+1}}$ are treated as negative passages $s^{-}$. Here, $m$ represents the number of hard negatives selected from $\mathcal{S}$. ColBERT produces two scores $R_{q,s^{+}}$ and $R_{q,s^{-}}$ from each triplet, and the model is optimized through a pairwise softmax cross-entropy loss:
(6) $\mathcal{L}_{ColBERT}=-\log(\frac{\exp R_{q,s^{+}}}{\exp R_{q,s^{+}}+\exp R_{q,s^{-}}})$
4. Experiments
4.1. Task and Datasets
To assess the effectiveness of W-RAG, we study three crucial components: the quality of the LLM-generated weak labels, the retrieval performance of the weakly trained dense retriever, and the overall effectiveness of the RAG system on the OpenQA task. We conduct experiments on four well-known datasets for OpenQA: MSMARCO QnA v2.1 (Nguyen et al., 2016), NQ (Kwiatkowski et al., 2019), SQuAD (Rajpurkar et al., 2016), and WebQ (Berant et al., 2013). MSMARCO QnA v2.1 shares the same corpus as MSMARCOv1 Passage Retrieval, with questions originating from real user queries submitted to Bing, making the language conversational. NQ also features real user questions and uses a corpus of documents from the English Wikipedia. In this work, we use a chuncked version of NQ as prepared by Karpukhin et al. (Karpukhin et al., 2020a). SQuAD’s corpus is similarly derived from the English Wikipedia, where the passages were first retrieved and then sampled. The questions and answers are manually written by crowdworkers, so they do not reflect natural language as closely as questions from real user queries like MSMARCO and NQ. WebQ’s corpus is drawn from Freebase, a large knowledge graph. Due to the nature of knowledge graphs, WebQ’s questions and answers are entity-related and factoid-based, thus making them less conversational compared to other datasets like MSMARCO and NQ.
For each dataset, we uniformly random sampled 5,000 question-answer pairs and a corpus of 500,000 passages from the training set, ensuring that all questions had relevant passages included in the corpus. We then split the question-answer pairs into 2,000 as training set, 1,000 as validation set, and 2,000 as test set. This sampling was necessitated by resource and time constraints. We argue that this small training set is sufficient to demonstrate statistically significant differences in both retrieval and the final OpenQA performance. While our results are indicative, our method could benefit from further validation using larger datasets to ensure generalizability.
4.1.1. Weak Label Quality
For our main experiments, we use Llama3-8B-Instruct (Dubey and .etl, 2024) to serve as the reranker. The 2,000 question-answer pairs in the training set are used to generate weak labels. For each question, weak labels are generated by scoring the top 100 passages retrieved by BM25, based on the LLM’s likelihood of generating the answer. This process requires the LLM to perform inference 100 times to produce a ranked list. The choice of 100 passages is a trade-off between accuracy and latency. We found that BM25’s Recall@100 reaches approximately 80% across all four datasets, and retrieving additional passages yields diminishing returns as the LLM’s inference time increases linearly. We evaluate the reranking performance using different prompts and LLMs, which will be elaborated on later in the ablation studies (5.4).
4.1.2. Weakly Trained Retriever
After the LLM weakly labels the training set, we use these labels to train DPR and ColBERT. For DPR, since in-batch negatives are used during training, we only need to select the top-ranked passage to curate a list of relevant question-passage pairs, thus forming 2,000 training samples. For all four datasets, we used two different initializaiton settings for DPR: one is using bert-base-uncased to train from scratch (denoted as “DPRbase”), and the other by resuming training from an unsupervised dense retriever using Yibin-Lei/ReContriever (Lei et al., 2023b) (denoted as “DPRReCon”).
ColBERT’s training process differs slightly from that of DPR because it requires the sampling of negative passages. For each question, we select the top-ranked passage from the reranked list as the positive passage and the immediate following 10 passages as hard negatives. These hard negatives are relevant but not sufficiently informative to elicit the correct answer from the LLM. We did not explore varying the number of hard negatives, as this is not the primary focus of this work. Additionally, since the off-the-shelf ColBERTv2 is trained on the entire MSMARCO dataset, using it as the starting point would introduce data leakage. For consistency, we always train ColBERT from scratch using bert-base-uncased for all four datasets (denoted as “ColBERTbase”).
4.1.3. OpenQA Performance
Once the dense retriever is trained, we integrate it into the generic RAG pipeline, where it is used to retrieve the top-$K$ evidence passages for a given user question. The retrieved passages are directly inserted into the prompt, as shown in Figure 2 alongside the question and instructions asking the LLM to generate an answer based on the retrieved passages or, if the passages are not useful, to answer using its internal knowledge. For our main experiments, we restrict Llama3-8B-Instruct to generate a maximum of 20 tokens and only use top 1 retrieved passage to supplement the prompt. We also explored the impact of using different numbers of supplemental evidence passages on both OpenQA performance and latency, which we discuss in the ablation studies (5.4). We did not experiment with different LLMs for the OpenQA task, as the LLM component of RAG is not the primary focus of our study.
4.2. Baselines
Since our retriever is trained from scratch with limited data, we do not compare it to state-of-the-art retrievers or RAG systems. Instead, we focus on the improvement from untrained baselines to our weakly trained models and compare them with models trained on ground-truth data.
For OpenQA evaluation, we use Llama3-8B-Instruct as the LLM in the RAG pipeline, assessing the impact of different retrievers on final OpenQA performance. We compare these retrievers against the same baseline models used for OpenQA. Additionally, we introduce two baselines: “Naive,” where Llama3 answers the question without any external information, showcasing its parameter knowledge, and “Groundtruth,” where the ground-truth passage is inserted into the prompt, representing the best possible OpenQA performance.
To assess weak label quality, we compare Llama3-8B-Instruct with BM25 after Llama3 generates weak labels on the training set of each dataset. For retriever performance, we evaluate two categories of models. The first category includes entirely unsupervised models: BM25, ColBERT initialized with bert-base-uncased (denoted as “ColBERTinit”), Contriever (Izacard et al., 2021), and ReContriever (Lei et al., 2023b). Contriever learns by contrasting sampled passages from different documents, assuming that each document is somewhat unique, while ReContriever enhances this by estimating relevance among sampled passages. The second category includes DPRReCon and ColBERTbase, both trained on ground-truth data. For a fair comparison, we use the same questions and an equal number of positive and negative samples as training data.
Retriever MSMARCO QnA NQ SQuAD WebQ
F1 Rouge-L BLEU-1 F1 Rouge-L BLEU-1 F1 Rouge-L BLEU-1 F1 Rouge-L BLEU-1
Naive 0.2749 0.2392 0.2019 0.1158 0.1136 0.0618 0.0885 0.0845 0.0508 0.1512 0.1487 0.0823
Groundtruth 0.4677 0.4310 0.3541 0.2186 0.2171 0.1193 0.2691 0.2675 0.1613 0.1736 0.1707 0.0937
ColBERTinit 0.2185 0.1867 0.1619 0.0525 0.0518 0.0283 0.0501 0.0482 0.0296 0.0434 0.0423 0.0233
BM25 0.3060 0.2712 0.2292 0.1374 0.1358 0.0748 0.1387 0.1369 0.0835 0.1085 0.1055 0.0594
Contriever 0.3125 0.2798 0.2347 0.1446 0.1432 0.0796 0.1144 0.1124 0.0680 0.1273 0.1257 0.0702
ReContriever 0.3171 0.2854 0.2387 0.1440 0.1429 0.0789 0.1257 0.1233 0.0753 0.1274 0.1239 0.0695
Trained on Groundtruth Data
ColBERTbase 0.3227 0.2876 0.2421 0.1494 0.1482 0.0814 0.1524 0.1497 0.0900 0.1006 0.1001 0.0538
DPRReCon 0.3424 0.3102 0.2588 0.1641 0.1628 0.0902 0.1550 0.1531 0.0919 0.1397 0.1378 0.0771
Trained on W-RAG Data
ColBERTbase 0.3150 0.2803 0.2365 0.1313 0.1301 0.0712 0.1509 0.1489 0.0897 0.1003 0.0989 0.0545
DPRReCon 0.3397 0.3074 0.2559 0.1528 0.1515 0.0837 0.1425 0.1402 0.0850 0.1304 0.1276 0.0727
Table 1. OpenQA Performance. Top 1 passage is inserted into the OpenQA prompt, limit number of generated tokens to 20. : Performance improvements are statistically significant with a ttest of $p<0.05$. DPRReCon’s baseline is ReContriever, ColBERTbase’s baseline is ColBERTinit
4.3. Experimental Settings
We begin by retrieving the top 100 passages using BM25 with the rank_bm25  [^3] https://github.com/dorianbrown/rank_bm25 BM25Okapi with $k_{1}$=1.5, $b$=0.75 and $\epsilon$=0.25. The nltk  [^4] https://github.com/nltk/nltk tokenizer is used to tokenize the questions and passages. We use BEIR  [^5] https://github.com/beir-cellar/beir to manage data, and the DPR models are trained using sentence-transformers with the following hyperparameters: batch size of 128, learning rate of $2\times 10^{-5}$, AdamW optimizer, 20 epoch, and model weights are saved based on the best Recall@5 on the validation set. ColBERT is trained from scratch using RAGatouille  [^6] https://github.com/bclavie/RAGatouille with a batch size of 64, 10 hard negatives per positive, learning rate of $1\times 10^{-5}$, AdamW optimizer, and 1 epoch. Generating weak labels took approximately two days per dataset on an Nvidia V100 GPU. Our experiments are easily reproducible since BM25, ColBERT, DPR, Llama3, and all four datasets are publicly accessable.
The prompts used to generate weak labels are shown in Figure 2. We use the term “DOCUMENT” instead of “PASSAGE” because our testing experiments showed that “DOCUMENT” consistently yields better reranking performance. We suspect this is because “DOCUMENT” is a more common term and likely appears more frequently in Llama3’s training data, leading it to a deeper understanding of the word “DOCUMENT”. Additionally, we tested different orderings of the passage, question, and instructions, finding that the Passage-Question-Instruction format consistently outperforms Passage-Instruction-Question and Instruction-Passage-Question.
4.4. Evaluation
We use Recall as our primary metric to evaluate the retrieval performance, as it focuses on whether the relevant passage is included within a specific ranking range. We also report MRR (Mean Reciprocal Rank), which is similar to Recall but also considers the position at which the relevant passage is ranked.
To gauge the alignment between machine-generated answers and ground-truth answers, we use F1, Rouge-L (Lin, 2004), and BLEU-1 (Papineni et al., 2002). These metrics collectively measure the overlap and sequence matching between generated and actual answers. We evaluate OpenQA performance on our uniformly random sampled test set, which contains 2,000 questions for each datasets. For questions with multiple correct answers in the dataset, we select the answer with the highest cumulative score across all OpenQA metrics.
Retriever MSMARCO QnA NQ SQuAD WebQ
MRR@5 R@1 R@5 MRR@5 R@1 R@5 MRR@5 R@1 R@5 MRR@5 R@1 R@5
ColBERTinit 0.0163 0.0120 0.0220 0.0866 0.0093 0.0224 0.0504 0.0063 0.0158 0.0031 0.0001 0.0004
BM25 0.2781 0.1647 0.4580 0.5686 0.0869 0.3575 0.2978 0.0400 0.2011 0.2737 0.0285 0.1336
Contriever 0.2826 0.1585 0.4901 0.5524 0.0988 0.3585 0.2221 0.0294 0.1606 0.2990 0.0165 0.1304
ReContriever 0.2908 0.1589 0.5093 0.5831 0.1118 0.3952 0.2150 0.0284 0.1597 0.3022 0.0194 0.1370
Trained on Groundtruth Data
$\text{DPR}_{base}$ 0.3237 0.2117 0.4994 0.5608 0.1174 0.3619 0.1786 0.0199 0.1054 0.4028 0.0561 0.2059
ColBERTbase 0.3484 0.2097 0.5112 0.5925 0.1316 0.3082 0.2310 0.0469 0.1072 0.1157 0.0180 0.0501
DPRReCon 0.3652 0.2284 0.5913 0.6632 0.1516 0.4715 0.2829 0.0393 0.2072 0.4155 0.0640 0.2262
Trained on W-RAG Data
$\text{DPR}_{base}$ 0.2981 0.1809 0.4813 0.5302 0.1012 0.3364 0.1781 0.0210 0.1020 0.3377 0.0478 0.1656
ColBERTbase 0.3116 0.1973 0.4904 0.5229 0.0982 0.2672 0.2265 0.0445 0.1029 0.0831 0.0153 0.0326
DPRReCon 0.3496 0.2023 0.5887 0.6301 0.1311 0.4267 0.2811 0.0382 0.2001 0.3610 0.0499 0.1880
Table 2. Retrieval Performance after training with W-RAG or ground-truth data. : Performance improvements are statistically significant with a ttest of $p<0.05$.
Retriever MSMARCO QnA NQ SQuAD WebQ
R@1 R@5 R@50 R@1 R@5 R@50 R@1 R@5 R@50 R@1 R@5 R@50
BM25 0.1694 0.4464 0.7577 0.088 0.353 0.7941 0.0449 0.2069 0.6461 0.0217 0.1155 0.5765
Llama3-8B 0.5239 0.6614 0.7517 0.1542 0.4224 0.7303 0.0805 0.3088 0.6774 0.0494 0.1887 0.5502
Table 3. BM25 retrieves the top 100 passages, which are then reranked by Llama3-8B. The reranked lists are used as weak labels for training a dense retriever.
5. Experimental Results
In this section we discuss the final OpenQA performance, retrieval performance of trained retrievers, the quality of weak labels, and ablation studies.
5.1. Main Results
We do not claim that W-RAG outperforms state-of-the-art RAG methods or achieves the best retrieval performance on any of the four datasets. Instead, we highlight the performance improvements of weakly trained retrievers compared to baselines, and demonstrate that the gap between weakly supervised and ground-truth trained retrievers is relatively small. This relationship is evident in both the retrieval performance and the final RAG for OpenQA performance.
The goal of W-RAG is to enhance LLM generation quality by providing relevant evidence passages. Tabel 1 presents the main OpenQA results using different retrievers in the RAG pipeline, evaluated on 2,000 test questions for each dataset. For these experiments, only the top 1 retrieved passage is added to the prompt, and the LLM is limited to generating 20 tokens. The Naive and Groundtruth baselines represent the lower and upper bounds, respectively. The Naive baseline excludes any supplementary passage, while Groundtruth includes the best passage. Any retriever that retrieves relevant passages should outperform the Naive baseline, while surpassing the Groundtruth baseline would be very unexpected. Results show consistent improvements with W-RAG trained retrievers across all datasets, with most of them being statistically significant. The relatively small gap between W-RAG trained and ground-truth trained retrievers suggests that W-RAG data approaches the quality of human-labeled data.
One baseline, ColBERTinit (initialized with bert-base-uncased), underperforms the Naive baseline, likely due to irrelevant passages being retrieved, introducing noise that misguides Llama3. We notice that for NQ, SQuAD, and WebQ, all the OpenQA metrics are small. This is due to the brevity of ground-truth answers (average length  2 words). For example, in response to the NQ question ”When did Big air Snowboarding become an Olympic sport?”, Llama3 outputs the correct answer: ”Big Air Snowboarding was first added to the 2018 Winter Olympics.” Since the ground-truth answer given by NQ is just ”2018”, standard OpenQA metrics like F1 and Rouge-L would yield low scores due to the denominator of the Precision term being the length of the generated answer, while the numerator is n-gram overlap which is at most 1 in this case. Despite this, these metrics remain valuable; if Llama3’s answer omitted ”2018,” the score would be zero. The absolute metric values are less important than the relative differences across settings. Although token length could be treated as a hyperparameter, we limited generation to 20 tokens across datasets for consistency.
5.2. Retrieval Results
The retrieval results of different trained retrievers are presented in Table 2. These results are based on 2,000 test questions for each trained or baseline retriever. On the MSMARCO QnA dataset, all retrievers trained on W-RAG data outperformed all baselines, while DPRReCon consistently outperformed all unsupervised baselines, except for SQuAD where BM25 performed best. As expected, DPR trained on ReContriever showed better retrieval performance than DPR trained from scratch. Similar to the OpenQA task, W-RAG trained retrievers slightly underperform compared to those trained on ground-truth data.
Although there appears to be a positive correlation between retrieval performance and OpenQA performance, this relationship is noisy and sometimes inconsistent. For instance, W-RAG trained ColBERTbase achieved nearly 4% higher Recall@1 than ReContriever on MSMARCO QnA, ranking the relevant passage at top for about 80 more questions. However, when these top 1 passages are used for OpenQA, ColBERTbase performed slighly worse than ReContriever. Similarly, ground-truth trained DPRReCon significantly outperformed W-RAG trained DPRReCon in Recall@1 on MSMARCO QnA, yet their OpenQA performance was almost identical when using the top-ranked passage.
These findings suggest that the traditional definition of relevance may not directly impact the answer quality in RAG systems. W-RAG’s top passage is selected based on its ability to elicit the correct answer from the LLM, while ground-truth passages are typically chosen based on term overlap, semantic similarity, or human judgment. Contrary to expectations, better retrieval does not always lead to better OpenQA. This raises the question: “Is the relevance definition in existing datasets truly effective for evaluating retrievers in RAG systems?”
5.3. W-RAG Labels
After generating all of W-RAG data, we examine its retrieval performance and the quality of the weakly labeled rank lists. In Table 3, a significant gap between BM25 and Llama3’s reranking performance at Recall@1 is evident, which justifies our choise of only using the top 1 passage as the positive when training different retrievers. We also observe that Llama3’s performance declines as the number of passages increases. This is likely due to the fact that starting from around the 50th position, the difference in answer likelihood between adjacent passages diminishes to negligible values, such as $\exp(-3.24815)-\exp(-3.24899)\approx 3\times 10^{-5}$.
Prompt R@1 R@5 F1 Rouge-L BLEU
BM25 0.1343 0.4128 0.2797 0.2436 0.0640
UPR 0.1970 0.4970 0.2898 0.2538 0.0677
Zero-shot 0.5027 0.6289 0.4300 0.3924 0.1876
One-shot 0.5187 0.6690 0.4358 0.3976 0.1939
Two-shot 0.5137 0.6840 0.4256 0.3858 0.1812
Table 4. Llama3-8B-Instruct’s reranking performance using different prompts on 500 validation questions from the MSMARCO QnA dataset. Answers are generated using only the top 1 reranked passage.
5.4. Ablation Study
We assessed the quality of W-RAG data generated by different LLMs, specifically their ability to rerank passages based on answer likelihood. This is illustrated in Figure 3, showing the reranking performance of various LLMs on 500 questions from the validation set of MSMARCO QnA. Although the LLMs notably differs at Recall@5, they exhibit consistent trends with each other. We notice Llama3-8B-Instruct is weak when compared to other examined LLMs. However, since their recall distributions are similar, any of these LLMs can serve as the reranker. We chose to use Llama3-8B-Instruct for all our experiments.
# F1 Rouge-L BLEU-1 Latency (sec/infer)
1 0.3397 0.3074 0.2559 1.043
3 0.3744 0.3395 0.2818 1.270
5 0.3767 0.3427 0.2835 1.454
10 0.3864 0.3524 0.2911 2.134
Table 5. Adding different # of evidence passages to the prompt when doing OpenQA using DPRReCon on 2,000 test questions from MSMARCO QnA.
Using Llama3-8B-Instruct, we evaluated reranking and OpenQA performance, as presented in Table 4. UPR (Sachan et al., 2022a), which ranks passages based on question likelihood, serve as an unsupervised approach for generating training labels. With weak supervision of ground-truth answers, W-RAG’s zero-shot, one-shot, and two-shot prompts significantly outperforms BM25 and UPR in both reranking and OpenQA metrics. However, the differences between the various shot prompts are not very consistent and are not significant. We hypothesize that this inconsistency may have stemmed from the LLM misinterpreting example passages as the relevant passage, or if the question types are drastically different between the example and the given question, thus introducing noise and reducing result reliability. Due to this, we conducted all subsequent experiments using zero-shot prompts, shown in Figure 2.
During the final OpenQA, we also study the impact of giving more evidence passages, shown in Table 5. We observe a consistent and steady increase in all OpenQA metrics when more evidence passages are inserted into the prompt. This behavior is expected because the more passages supplied, the better chance a good passage is within them. However, latency also grows as we increase the number of passages.
Figure 3. Comparison of recall for various LLMs at different top k positions, when reranking top 100 passages retrieved by BM25.
6. Conclusions and Future Work
This paper introduces a general framework, W-RAG, for extracting weak labels from question-answer pairs to address the scarcity of training data for dense retrieval in RAG for OpenQA. W-RAG reranks the top-$K$ initially retrieved passages by the probability of LLMs generating the OpenQA answer conditioned on each question-passage pair. This ranking score computation method aligns well with the latest study (Cuconasu et al., 2024), which states that the retrieved passage should answer the question; otherwise, performance will be negatively impacted. Extensive experimental results on four public OpenQA datasets demonstrate the effectiveness of W-RAG.
In future work, we plan to explore which types of passages most effectively enhance RAG’s performance in OpenQA, as indicated by (Cuconasu et al., 2024), where even randomly sampled tokens were beneficial in some cases. Understanding the types of passages preferred by LLMs will allow us to design more effective dense retrieval methods, including new structures and evaluation metrics. Additionally, the compression of retrieved passages warrants further study, as directly feeding all retrieved passages to LLMs not only increases computational complexity but also introduces significant noise. Finally, even with a ground-truth evidence passage, RAG can still produce incorrect answers, known as hallucinations; enhancing the robustness of RAG in OpenQA is another promising direction for future research.

### v2

@@
 Abstract.
-In knowledge-intensive tasks such as open-domain question answering (OpenQA), Large Language Models (LLMs) often struggle to generate factual answers relying solely on their internal (parametric) knowledge.
+In knowledge-intensive tasks such as open-domain question answering (OpenQA), large language models (LLMs) often struggle to generate factual answers, relying solely on their internal (parametric) knowledge.
 To address this limitation, Retrieval-Augmented Generation (RAG) systems enhance LLMs by retrieving relevant information from external sources, thereby positioning the retriever as a pivotal component.
 Although dense retrieval demonstrates state-of-the-art performance, its training poses challenges due to the scarcity of ground-truth evidence, largely attributed to the high costs of human annotation.
-In this paper, we propose W-RAG by utilizing the ranking capabilities of LLMs to create weakly labeled data for training dense retrievers.
-Specifically, we rerank the top-$K$ passages retrieved via BM25 by assessing the probability that LLMs will generate the correct answer based on the question and each passage.
-The highest-ranking passages are then used as positive training examples for dense retrieval.
-Our comprehensive experiments across four publicly available OpenQA datasets demonstrate that our approach enhances both retrieval and OpenQA performance compared to baseline models.
-Source code is published  [^1] https://github.com/jmnian/weak_label_for_rag.
-Retrieval Augmented Generation, Open-domain Question Answering, Weak Supervised Learning, Dense Retrieval
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
+In this paper, we propose W-RAG, a method that draws weak training signals from the downstream task (such as OpenQA) of an LLM, and fine-tunes the retriever to prioritize passages that most benefit the task.
+Specifically, we rerank the top-$k$ passages retrieved via BM25 by assessing the probability that the LLM will generate the correct answer for a question given each passage.
+The highest-ranking passages are then used as positive fine-tuning examples for dense retrieval.
+We conduct comprehensive experiments across four publicly available OpenQA datasets to demonstrate that our approach enhances both retrieval and OpenQA performance compared to baseline models, achieving results comparable to models fine-tuned with human-labeled data.
+Source code is published  [^1] https://github.com/jmnian/WRAG.
 1. Introduction
@@
 Previous studies (Lewis et al., 2020; Xiong et al., 2020a) adopt a seq2seq model as a reader and train it on the labeled dataset.
-Recently, large language models (LLMs) like GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023) have demonstrated astonishing performance on various tasks, including OpenQA, attributing to the substantial amount of knowledge stored in their internal parameters.
+Recently, large language models (LLMs) like GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023) have demonstrated astonishing performance on various tasks, attributed to the substantial amount of knowledge stored in their internal parameters.
 Despite the unprecedented achievements of LLMs, they face constraints such as the inability to consistently integrate up-to-date knowledge, as their parametric knowledge is fixed after being trained on huge datasets.
@@
 Neither cross-attention scores nor the likelihood of LLMs generating the question based on the input passage explicitly ensures that the question can be answered by the “positive” passage.
-To address the scarcity of training data for dense retrievers in RAG for OpenQA, we propose extracting weak labels from existing OpenQA question-answer pairs by leveraging the ranking capabilities of LLMs (Sun et al., 2023).
+To address the scarcity of training data for dense retrievers in RAG for OpenQA, and to bridge the gap between semantic relevance and containing the actual answer, we propose extracting weak labels from existing OpenQA question-answer pairs by leveraging the ranking capabilities of LLMs (Sun et al., 2023).
 Specifically, we first use BM25 to retrieve the top-$k$ passages for a question, then pair each passage with the question.
@@
 Only the top-ranked passage is selected as the positive example for the question, and we train the dense retrievers using in-batch negative sampling.
-Our method evaluates the relevance of the question-passage pair by the likelihood of the answer given a passage and the question, making it particularly suitable for OpenQA, where retrieved passages must be capable of providing the correct answer.
+Our method evaluates the relevance of a question-passage pair by the likelihood of generating the correct answer for the question given the passage, making it particularly suitable for OpenQA, where retrieved passages must be capable of providing the correct answer.
 We conducted comprehensive experiments across four publicly available OpenQA datasets, and the results demonstrate that our approach enhances both retrieval and OpenQA performances compared to baseline models.
 Our contributions can be summarized as:
-  • We propose W-RAG, a general framework to generate weak labels from question-answer pairs for training dense retrievers in RAG.
-  • Comprehensive experiments were conducted, and the results prove that W-RAG improves both retrieval and OpenQA performance over baselines.
-  • Open source LLMs were used, and the code was released anonymously to ensure reproducibility.
+  • We propose W-RAG, a general framework that uses answer likelihoods from an LLM given question-answer pairs to generate weak relevance labels for fine-tuning dense retrievers in RAG systems.
+  • Comprehensive experiments demonstrate that W-RAG data improves both retrieval and OpenQA performance over baselines and achieves results comparable to models trained on human-labeled data.
+  • Open source LLMs were used, and the code was released to ensure reproducibility.
 2. Related Work
 2.1. Dense Retrieval
-Traditional information retrieval (IR) method are based on exact term matching, like BM25.
+Traditional information retrieval (IR) methods are based on exact term matching, like BM25.
 While it is still widely used due to its efficiency, effectiveness, and robustness, it suffers from the well-known issue of lexical gap (Berger et al., 2000).
-To address this, leveraging the neural networks, dense retrieval (DR) employs pre-trained language models like BERT (Devlin et al., 2019), to encode questions and passages into embeddings, and measures the question-passage relevance score by the vector similarly in the embedding space.
-Specifically, DP encodes the whole corpus into embeddings and builds the index, such as Faiss (Douze et al., 2024), on them.
+To address this, leveraging neural networks, dense retrieval (DR) employs pre-trained language models like BERT (Devlin et al., 2019) to encode questions and passages into embeddings, and measures the question-passage relevance score by comparing the vectors similarly in the embedding space.
+Specifically, DR encodes the whole corpus into embeddings and builds the index, such as Faiss (Douze et al., 2024), on them.
 When a new question comes in, DR encodes the question into an embedding and performs a nearest neighbor search.
 DR can be classified into two categories: supervised, like DPR (Karpukhin et al., 2020b), TAS-B (Hofstätter et al., 2021), and ColBERT (Khattab and Zaharia, 2020a; Santhanam et al., 2022); unsupervised, like Contriever(Izacard et al., 2022) and ReContriever(Lei et al., 2023a), based on the training method.
-DPR utilizes a dual-tower architecture, with one BERT model dedicated to encode questions and another to encode passages.
-The similarity between the question and passage embeddings is then calculated, aiming to maximize the log-likelihood of the positive passage.
+DPR utilizes a dual-tower architecture, with one BERT model dedicated to encoding questions and another to encoding passages, though in practice sometimes a single encoder is shared between passages and questions.
+The similarity between the question and passage embeddings is calculated, aiming to maximize the log-likelihood of the positive passage.
 ColBERT uses the same BERT model for both the question and passage encoders, differentiating them by appending a unique special token after the [CLS] token.
@@
 RAG models have been applied to OpenQA, demonstrating significant performance improvement.
-Different RAG models have been proposed to solve the critical issues in OpenQA, such as how to retrieve relevant passages (Yu et al., 2023; Shi et al., 2023; Shao et al., 2023; Trivedi et al., 2023), when to call the retriever (Jiang et al., 2023b; Wang et al., 2023; Cheng et al., 2024), and how to decrease the computational complexity (Li et al., 2023; Jiang et al., 2023a; Xu et al., 2024; Kim et al., 2024).
-Since the retriever is a critical component in RAG, some studies have tried to improve the quality of retrieved passages, including training a better retriever (Yu et al., 2023; Shi et al., 2023) and prompt engineering (Shao et al., 2023; Trivedi et al., 2023).
-REPLUG (Shi et al., 2023) trains a dense retriever by minimizing the KL divergence between the retrieval likelihood and the language model likelihood computed as the normalized language model probability of the ground-truth, given the question and passage.
-AAR (Yu et al., 2023) combines the top-$K$ passages ranked by the LLM’s averaged corss-attention scores with ground-truth passages as positive passages and then follow ANCE (Xiong et al., 2021) to sample negative passages to train the dense retriever.
+Different RAG models have been proposed to solve the critical issues in OpenQA, such as how to retrieve relevant passages (Yu et al., 2023; Shi et al., 2024; Shao et al., 2023; Trivedi et al., 2023), when to call the retriever (Jiang et al., 2023c; Wang et al., 2023; Cheng et al., 2024), and how to decrease the computational complexity (Li et al., 2023; Jiang et al., 2023b; Xu et al., 2024; Kim et al., 2024).
+Since the retriever is a critical component in RAG, some studies have tried to improve the quality of retrieved passages, including training a better retriever (Yu et al., 2023; Shi et al., 2024) and prompt engineering (Shao et al., 2023; Trivedi et al., 2023).
+AAR (Yu et al., 2023) combines the top-$k$ passages ranked by the LLM’s averaged cross-attention scores with ground-truth passages as positive passages and then follows ANCE (Xiong et al., 2021) to sample negative passages to train the dense retriever.
 ITER-RETGEN (Shao et al., 2023) leverages the model output from the previous iteration as a specific context to help retrieve more relevant knowledge.
-IRCoT (Trivedi et al., 2023) also adopt a similar approach to perform retrieval, but applies CoT for generating responses.
+IRCoT (Trivedi et al., 2023) also adopts a similar approach to perform retrieval, but applies CoT for generating responses.
 In some cases, LLMs can generate factual content well without external knowledge, and retrieving passages will decrease RAG’s performance.
-Self-RAG (Wang et al., 2023) and FLARE (Jiang et al., 2023b) both fine-tune the LLMs to call the search engine automatically when external knowledge is needed.
+Self-RAG (Wang et al., 2023) and FLARE (Jiang et al., 2023c) both fine-tune the LLMs to call the search engine automatically when external knowledge is needed.
 UAR (Cheng et al., 2024) trains classifiers to identify the need for external knowledge.
@@
 Selective-Context (Li et al., 2023) filters non-essential lexical units by the summarization of self-information of each token contained in the unit.
-LongLLMLingua (Jiang et al., 2023a) contrasts the perplexity score of each token in the passage with the perplexity score of the same token conditioned on the question and adopts this conservative perplexity score to filter out tokens.
+LongLLMLingua (Jiang et al., 2023b) contrasts the perplexity score of each token in the passage with the perplexity score of the same token conditioned on the question and adopts this conservative perplexity score to filter out tokens.
 RECOMP (Xu et al., 2024) leverages the summarization models to summarize the retrieved passages.
-SuRe (Kim et al., 2024) summarizes retrieved passages conditioned on each answer candidates generated by prompting the LLMs and then selects the top summarizations ranked by LLMs through a combination of pointwise and pairwise scoring method.
+SuRe (Kim et al., 2024) summarizes retrieved passages conditioned on each answer candidate generated by prompting the LLMs and then selects the top summarizations ranked by LLMs through a combination of pointwise and pairwise scoring methods.
 Figure 1.
-W-RAG fits into the general RAG pipeline by training the retriever with LLM generated weak labels.
+W-RAG fits into the general RAG pipeline by training the retriever with LLM-generated weak labels.
 Following the steps from top to bottom, we retrieve candidate passages using BM25, present each passage to the LLM to rerank based on the answer likelihood, then use the reranked top passage to train the retriever to finally enhance LLM’s response quality through the standard RAG pipeline.
-The entire prompt is shown in Figure 2.
+Prompt in “Step 2” is shown in Figure 2.
+2.3. LLM Signals for Retriever Training
+REPLUG (Shi et al., 2024) improves dense retrieval by minimizing the KL divergence between the retrieval likelihood and LM likelihood of generating the second half of a text given the first half and a retrieved passage.
+The higher the likelihood that the LM will generate the second half correctly, the more relevant the retrieved passage is considered.
+Since only two separate corpora of text are needed to mine training signals, the method is highly scalable.
+CoRAG (Wang et al., 2025) while focusing on multi-hop retrieval, selects training samples based on retrieval chains that maximize answer likelihood from the LLM.
+(Li et al., 2024) propose a teacher-student distillation method where they prompt the LLM to directly output ranked lists to fine-tune a reranker via list-wise ranking loss, while simultaneously aligning the retriever’s output distribution with the reranker’s.
+Similarly, ARL2 (Zhang et al., 2024) scores relevance by prompting LLMs to annotate passages as either supporting or not supporting the answer for a given question, then trains the retriever to align with these annotations.
 2.4. LLMs for Ranking
-Concerning the generation of weakly labeled ranking data using LLMs as passage rankers, existing work can be categorized into three categories.
-The first involves a zero-shot listwise ranking strategy, where a prompt containing the question, instructions, and multiple passages are given to the LLM, and the LLM generates a permutation of ranked indices (Baldelli et al., 2024; Ma et al., 2023; Pradeep et al., 2023; Shen et al., 2023).
-The second category prompts the LLM with passages and directly instructs it to rate their relevance (Zhuang et al., 2024), or perform pairwise comparisons (Qin et al., 2023).
+Concerning the generation of weakly labeled ranking data using LLMs as passage rankers, existing work can be broadly grouped into three categories.
+The first category involves a zero-shot listwise ranking strategy, where a prompt containing the question, instructions, and multiple passages is given to the LLM, and the LLM generates a permutation of ranked indices (Baldelli et al., 2024; Ma et al., 2023; Pradeep et al., 2023; Shen et al., 2023).
+The second category focuses on pointwise and pairwise relevance assessment.
+(Zhuang et al., 2024) prompts the LLM with a passage and directly instructs it to rate its relevance, making it a point-wise relevance assessment method.
+(Qin et al., 2023) performs pair-wise ranking where two passages are given in the prompt and the LLM is asked to point out which is more relevant to the question.
 The third approach calculates the question generation likelihood when a passage is included in the prompt (Cho et al., 2023; Sachan et al., 2022b; Zhuang et al., 2023).
-Our method for scoring passages aligns with this third approach; however, as mentioned in Section 1, we rank the passages by the likelihood of the LLM generating the question’s ground-truth answer based on the question-passage pair.
+Our method for scoring passages aligns with this third approach.
+However, as mentioned in the introduction, our method ranks passages based on the likelihood that the LLM generates the ground-truth answer when conditioned on the question-passage pair.
+The ranked lists are then used as weakly labeled training data to fine-tune the retriever, with the goal of teaching it to retrieve passages that are most likely to condition the LLM to generate a correct answer.
+Prompt for Answer Likelihood Extraction Passage: {passage} Question: {question} Please answer the question using the facts of the passage.
+Keep your answer grounded to the facts of the passage.
+Keep your answer within one short sentence.
+Answer: {answer}
+Figure 2. Bolded portion is evaluated for answer likelihood, which serves as a weak relevance score between “passage” and “question”.
 3. Method
-As illustrated in Figure 1, W-RAG introduces a novel approach for training the retrieval component within the RAG pipeline from scratch.
-Our method leverages a weakly-supervised training paradigm that requires only a set of question-answer pairs and an evidence corpus.
-The W-RAG process unfolds in three stages: first, it retrieves relevant passages from the evidence corpus; second, it employs a LLM to generate weak labels from them by reranking the retrieved passages by the likelihood of generating the ground-truth answer; finally, the dense retiever is trained using these weak labels.
-The resulting dense retriever is then capable of retrieving evidence, thereby enhancing the performance of LLMs across a wide array of tasks.
-Figure 2. Prompts used for weak label generation and question answering
+As illustrated in Figure 1, W-RAG introduces a novel approach for fine-tuning the retrieval component within the RAG pipeline from a pre-trained dense retriever.
+Our method employs a weakly supervised training paradigm that requires only a set of question-answer pairs and a passage corpus.
+The W-RAG process unfolds in three stages: First, a first-stage retriever, such as BM25, retrieves candidate passages from the corpus.
+Second, we use an LLM to estimate the relevance of each passage by computing the likelihood of generating the correct answer when conditioned on the given question and passage, producing a weakly labeled ranking list.
+Finally, the dense retriever is fine-tuned using these weak labels.
+The resulting dense retriever is optimized for ranking passages with higher answer likelihood towards top positions, thereby enhancing the performance of LLMs across a wide array of tasks.
 3.1. Weak-label Generation
-Given an evidence corpus $\mathcal{C}=\{d_{1},\dots,d_{N}\}$, we use a simple retreiver to conduct first-stage retrieval.
-Given a question $q$, the retriever ranks each passage by relevance and selects the top-$K$ passages $\mathcal{P}=\{s_{1},\dots,s_{K}\}$, where $\mathcal{P}\subset\mathcal{C}$.
-The corpus may comprise chuncked Wikipedia articles or from online sources such as Reddit and Common Crawl, while the question-answer pairs could be sourced from OpenQA platforms such as Quora and StackOverflow.
-The simple first-stage retriever could be lexical such as BM25 (Robertson and Zaragoza, 2009), or a dense retriever.
+Given a passage corpus $\mathcal{C}=\{s_{1},\dots,s_{N}\}$, we use a simple retriever to conduct first-stage retrieval.
+Given a question $q$, the retriever ranks each passage by relevance and selects the top-$k$ passages $\mathcal{P}=\{s^{\prime}_{1},\dots,s^{\prime}_{K}\}$, where $\mathcal{P}\subset\mathcal{C}$.
+The passage corpus may comprise chunked Wikipedia articles or those from online sources, such as Reddit and Common Crawl, while the question-answer pairs could be sourced from QA datasets and question-answering platforms, such as Quora and Stack Overflow.
+The simple first-stage retriever could be lexical, such as BM25 (Robertson and Zaragoza, 2009), or a dense retriever.
 The primary objective at this stage is to maximize recall within a manageable set of retrieved passages.
-Ideally, this step curates evidence passages containing the correct answer to the question
-W-RAG is motivated by the hypothesis that an autoregressive LLM is more likely to generate the correct answer if the provided passage contains the necessary information to answer the question.
-Our approach aims to harness signals from the LLM’s downstream tasks to evaluate the effectiveness of a passage in eliciting the correct answer.
-It is important to note that traditional relevance measures, such as term overlap or semantic similarity, do no necessarily guarantee that the most relevant passage contains the needed information.
-Instead, our target is on identifying passages that most effectively prompt the LLM to produce the correct answer and using these passages to train the dense retriever.
-In this context, we redefine passage relevance as the degree to which a passage can elicit the ground-truth answer, and this criterion serves as the target for training the dense retriever.
-As depicted in Figure 1 Step 2, we construct weak label generation prompts with candidate passage $s_{i}$, question $q$, some instructions $I$, and ground-truth answer $a$.
-The downstream signal is captured through the conditional probability of the ground-truth answer, denoted as $p(a|s_{i},q,I)$:
-(1) $p(a|s_{i},q,I)=\prod_{j}\text{LLM}(a_{j}|s_{i},q,I,a_{<j})$
-For each token in the input prompt, the LLM assigns logits, which after applying a softmax operation, correspond to the probability of each vocabulary token being the next in the sequence.
-The probability of the $j$th answer token, denoted as $\text{LLM}(a_{j}|s_{i},q,I,a_{<j})$, is directly extracted from the logits of the preceding token.
+Ideally, this step fetches evidence passages containing the correct answer to the question.
+W-RAG is based on the hypothesis that an autoregressive LLM is more likely to generate the correct answer when the input passage contains the necessary information.
+We aim to leverage this signal by fine-tuning the retriever to rank passages according to their likelihood of eliciting the correct answer.
+Unlike traditional relevance measures, such as term overlap or semantic similarity, which do not emphasize that a passage enables answer generation.
+Our approach redefines relevance as the extent to which a passage conditions the LLM to produce the ground-truth answer.
+These labels serve as a weak supervision signal for fine-tuning the retriever.
+As depicted in Figure 1 Step 2, we construct weak label generation prompts with candidate passage $s^{\prime}_{i}$, question $q$, some instructions $I$, and ground-truth answer $a$.
+The downstream signal is captured through the conditional probability of the ground-truth answer, denoted as $p(a|s^{\prime}_{i},q,I)$:
+(1) $p(a|s^{\prime}_{i},q,I)=\prod_{j}\text{LLM}(a_{j}|s^{\prime}_{i},q,I,a_{<j})$
+For each token in the input prompt, the LLM assigns logits, which, after applying a softmax operation, correspond to the probability of each vocabulary token being the next in the sequence.
+The probability of the $j$th answer token, denoted as $\text{LLM}(a_{j}|s^{\prime}_{i},q,I,a_{<j})$, is directly extracted from the logits of the preceding token.
 Since the ground-truth answer typically consists of multiple tokens, the cumulative probability can diminish rapidly.
-To address this, we simply take the logarithm on both sides of Equation 1, converting each token’s probability into log-likelihood.
+To address this, we take the logarithm on both sides of Equation 1, converting each token’s probability into log-likelihood.
 The overall probability of the answer $a$ is then computed as the average of the sum of the log-likelihoods for each token in the answer:
-(2) $\log p(a|s_{i},q,I)=\frac{1}{|a|}\sum_{j}\log\text{LLM}(a_{j}|s_{i},q,I,a_{<j})$
-We consider this log-likelihood as the relevance score for the candidate passage $s_{i}$.
-Each of the $K$ candidate passages is weakly labeled in this manner, resulting in a ranked list of passages, denoted as $\mathcal{S}=\{s_{r_{1}},\dots,s_{r_{K}}\}$.
+(2) $\log p(a|s^{\prime}_{i},q,I)=\frac{1}{|a|}\sum_{j}\log\text{LLM}(a_{j}|s^{\prime}_{i},q,I,a_{<j})$
+We consider this log-likelihood as the relevance score for the candidate passage $s^{\prime}_{i}$.
+Each of the $k$ candidate passages is weakly labeled in this manner, resulting in a ranked list of passages, denoted as $\mathcal{S}=\{s^{\prime}_{r_{1}},\dots,s^{\prime}_{r_{K}}\}$.
 This list is produced by sorting the set of candidate passages $\mathcal{S}$ according to their relevance scores.
@@
 3.2.1. DPR
-DPR utilizes a bi-encoder architecture, where the question and passage are independently mapped to an embedding space through two separate BERT encoders.
-For a given question $q$, the relevance score $R_{q,s}$ for each passage $s$ is given by the similarity between their respective [CLS] token embeddings, $\mathbf{e}_{q}$ and $\mathbf{e}_{s}$.
+DPR utilizes a bi-encoder architecture, where the question and passage are independently mapped into a shared embedding space through two separate BERT encoders.
+In most practical implementations, a single shared BERT encoder is used for both the question and the passage, which is also the configuration used in our experiments.
+For a given question $q$ and passage $s$, the relevance score $R_{q,s}$ is given by the similarity between their respective [CLS] token embeddings, denoted as $\mathbf{e}_{q}\in\mathbb{R}^{d}$ and $\mathbf{e}_{s}\in\mathbb{R}^{d}$.
 We use cosine similarity as the similarity function:
@@
 These pairs are then grouped into batches of size $n$ to form training batches, optimizing the retriever using a Multiple Negatives Ranking (MNR) loss (Henderson et al., 2017).
-In this approach, for the $i$th question-passage pair, $s_{i}$ is treated as the positive example, while the other passages within the same batch are treated as negative examples.
-The loss function is thus:
+In this approach, for the $i$th question-passage pair, $s_{i}$ is treated as the positive example, while the other passages within the same batch $\{s_{j}\}_{j\neq i}$ are treated as negative examples.
+The MNR loss function is thus:
 (4) $\mathcal{L}_{MNR}=-\sum_{i=1}^{n}\log(\frac{\exp(\alpha\cdot R_{q_{i},s_{i}})}{\sum_{j=1}^{n}\exp(\alpha\cdot R_{q_{i},s_{j}})})$
-The scaler $\alpha$ is used to amplify the cosine similarity score, usually set at 20 according to the default setting in sentence-transformers  [^2] https://sbert.net/docs/package_reference/sentence_transformer/losses.html#multiplenegativesrankingloss.
+The scaler $\alpha$ amplifies the cosine similarity score, typically set to 20 according to the default setting in sentence-transformers  [^2] https://sbert.net/docs/package_reference/sentence_transformer/losses.html#multiplenegativesrankingloss.
 3.2.2. ColBERT
-ColBERT employes a bi-encoder architecture with late interaction, where the question and passage are independently encoded using a shared BERT model to obtain embeddings $E_{q}$ and $E_{s}$.
-The relevance score for each passage is computed on the question side using a “MaxSim” operation, which sums the maximum similarities between any passage token embedding with each question token embedding:
-(5) $R_{q,s}=\sum_{i\in|E_{q}|}\max_{j\in|E_{s}|}\text{dot}(E_{q_{i}},E_{s_{j}})$
-Here, $\text{dot}(x,y)$ represents the dot product between $x$ and $y$.
+ColBERT employs a bi-encoder architecture with late interaction, where the question and passage are independently encoded using a shared BERT model to obtain embeddings $E_{q}\in\mathbb{R}^{l_{q}\times d}$ and $E_{s}\in\mathbb{R}^{l_{s}\times d}$, where $l_{q}$ and $l_{s}$ are the number of tokens in the question and passage, respectively, and $d$ is the embedding dimension.
+To compute the relevance score between a question and a passage, ColBERT applies a “MaxSim” operation: for each question token embedding $E_{q_{i}}\in E_{q}$, it finds the passage token embedding $E_{s_{j}}\in E_{s}$ that produces the highest dot product with $E_{q_{i}}$, and then sums these maximum values across all question tokens:
+(5) $R_{q,s}=\sum_{i=1}^{l_{q}}\max_{j\in[1,l_{s}]}\text{dot}(E_{q_{i}},E_{s_{j}})$
+Here, $\text{dot}(x,y)$ represents the dot product between vectors $x$ and $y$.
 The training dataset for ColBERT consists of triplet samples $\langle q,s^{+},s^{-}\rangle$, extracted from each ranked list $\mathcal{S}$.
-In these triplets, $s_{r_{1}}$ is treated as the positive passage $s^{+}$, while $s_{r_{2}},s_{r_{3}},\dots,s_{r_{m+1}}$ are treated as negative passages $s^{-}$.
+In each triplet, $s_{r_{1}}$ is treated as the positive passage $s^{+}$, while $s_{r_{2}},s_{r_{3}},\dots,s_{r_{m+1}}$ are treated as negative passages $s^{-}$.
 Here, $m$ represents the number of hard negatives selected from $\mathcal{S}$.
-ColBERT produces two scores $R_{q,s^{+}}$ and $R_{q,s^{-}}$ from each triplet, and the model is optimized through a pairwise softmax cross-entropy loss:
+ColBERT computes relevance scores $R_{q,s^{+}}$ and $R_{q,s^{-}}$ from each triplet, and optimizes the model using a pairwise softmax cross-entropy loss:
 (6) $\mathcal{L}_{ColBERT}=-\log(\frac{\exp R_{q,s^{+}}}{\exp R_{q,s^{+}}+\exp R_{q,s^{-}}})$
+Retriever NQ SQuAD WebQ MSMARCO QnA
+F1 EM BLEU-1 F1 EM BLEU-1 F1 EM BLEU-1 F1 EM BLEU-1
+Lower and upper bound of OpenQA performance
+Naive 10.14 0.15 5.64 22.40 11.62 19.99 38.96 17.30 31.84 26.92 0.25 20.19
+Ground truth 69.95 54.90 65.43 77.69 58.10 72.99 52.39 33.76 46.08 43.40 9.95 28.44
+Without fine-tuning
+BM25 52.32 38.40 48.08 47.01 32.90 43.62 36.63 20.68 31.38 30.53 5.16 19.81
+Contriever 54.84 41.05 50.54 41.28 28.85 38.22 39.28 19.62 33.16 30.84 5.28 19.79
+ColBERTinit 17.95 11.60 15.91 16.82 10.90 15.27 20.94 12.24 18.13 19.24 1.80 12.79
+ReContriever 55.01 40.75 50.64 43.50 30.10 40.19 37.79 20.89 31.74 31.11 5.42 20.06
+Fine-tuned with ground-truth data
+ColBERTbase 53.95 39.85 49.65 52.43 37.37 48.77 34.21 19.20 29.35 30.51 4.60 19.79
+DPRReCon 58.53 43.60 53.90 50.36 35.05 46.57 42.31 23.42 35.91 31.64 5.32 20.53
+Fine-tuned with W-RAG data
+ColBERTbase 50.35 37.75 46.42 52.28 36.55 48.32 32.16 18.99 27.69 29.79 4.95 19.01
+DPRReCon 56.40 42.00 51.98 49.71 34.95 45.99 41.59 23.63 35.30 32.41 5.23 20.85
+Table 1.
+OpenQA performance using Llama3.1-8B-Instruct.
+The top 5 passages retrieved by each retriever are inserted into the QA prompt.
+Bold indicates the best performance; Underlined indicates the second best (excluding “Naive” and “Ground truth”): Performance improvements are statistically significant with a t-test of $p<0.05$.
+Prompt can be found in Figure 3.
 4. Experiments
 4.1. Task and Datasets
-To assess the effectiveness of W-RAG, we study three crucial components: the quality of the LLM-generated weak labels, the retrieval performance of the weakly trained dense retriever, and the overall effectiveness of the RAG system on the OpenQA task.
+To assess the effectiveness of W-RAG generated weak labels, we study three crucial components: (1) the quality of the LLM-generated weak labels, (2) the retrieval performance of the weakly trained dense retriever, and (3) the end-to-end performance of the RAG system on the OpenQA task.
 We conduct experiments on four well-known datasets for OpenQA: MSMARCO QnA v2.1 (Nguyen et al., 2016), NQ (Kwiatkowski et al., 2019), SQuAD (Rajpurkar et al., 2016), and WebQ (Berant et al., 2013).
-MSMARCO QnA v2.1 shares the same corpus as MSMARCOv1 Passage Retrieval, with questions originating from real user queries submitted to Bing, making the language conversational.
+MSMARCO QnA v2.1 shares the same corpus as MSMARCOv1 Passage Retrieval, with questions originating from real user queries submitted to Bing.
+The language is generally conversational, with answers that are typically longer.
 NQ also features real user questions and uses a corpus of documents from the English Wikipedia.
-In this work, we use a chuncked version of NQ as prepared by Karpukhin et al.
+The answers are often very short and clean.
+In this work, we use a chunked version of the NQ corpus as prepared by Karpukhin et al.
 (Karpukhin et al., 2020a).
 SQuAD’s corpus is similarly derived from the English Wikipedia, where the passages were first retrieved and then sampled.
-The questions and answers are manually written by crowdworkers, so they do not reflect natural language as closely as questions from real user queries like MSMARCO and NQ.
+The questions and answers are manually written by crowdworkers, so they do not reflect natural language as closely as those from real user queries, such as MSMARCO QnA and NQ, and the answers are typically concise and focused.
 WebQ’s corpus is drawn from Freebase, a large knowledge graph.
-Due to the nature of knowledge graphs, WebQ’s questions and answers are entity-related and factoid-based, thus making them less conversational compared to other datasets like MSMARCO and NQ.
+Due to the nature of knowledge graphs, WebQ’s questions and answers are entity-related and factoid-based, making them less conversational.
+Similar to NQ and SQuAD, the answers in WebQ are generally short.
 For each dataset, we uniformly random sampled 5,000 question-answer pairs and a corpus of 500,000 passages from the training set, ensuring that all questions had relevant passages included in the corpus.
@@
 This sampling was necessitated by resource and time constraints.
-We argue that this small training set is sufficient to demonstrate statistically significant differences in both retrieval and the final OpenQA performance.
-While our results are indicative, our method could benefit from further validation using larger datasets to ensure generalizability.
+We argue that this small training set is sufficient to demonstrate statistically significant improvements in both retrieval and the final OpenQA performance.
+While our results are indicative, our method could benefit from further validation using more weakly labeled training data to ensure generalizability.
 4.1.1. Weak Label Quality
-For our main experiments, we use Llama3-8B-Instruct (Dubey and .etl, 2024) to serve as the reranker.
+For our main experiments, we use Llama3-8B-Instruct (Dubey and .etl, 2024) to serve as the weak supervision labeler.
 The 2,000 question-answer pairs in the training set are used to generate weak labels.
-For each question, weak labels are generated by scoring the top 100 passages retrieved by BM25, based on the LLM’s likelihood of generating the answer.
-This process requires the LLM to perform inference 100 times to produce a ranked list.
+For each question, we score the top 100 passages retrieved by BM25 based on the LLM’s likelihood of generating the ground-truth answer.
+We chose BM25 for its effectiveness and efficiency, though in principle this first-stage retrieval can be done with any retriever.
+The labeling process requires the LLM to perform 100 inference calls using the prompt shown in Figure 2 to produce a ranked list.
 The choice of 100 passages is a trade-off between accuracy and latency.
 We found that BM25’s Recall@100 reaches approximately 80% across all four datasets, and retrieving additional passages yields diminishing returns as the LLM’s inference time increases linearly.
-We evaluate the reranking performance using different prompts and LLMs, which will be elaborated on later in the ablation studies (5.4).
+We evaluate the reranking performance using different LLMs, which will be elaborated on later in the ablation studies (5.4).
+Question Answering Prompt [System Prompt] You are a helpful assistant who answers questions using the provided passage.
+If the passage does not contain the necessary information, you will answer the question directly based on your general knowledge.
+Only provide the final answer, no explanations.
+[User Prompt] Passage: {passage #1} Passage: {passage #2} Passage: {passage #3} Passage: {passage #4} Passage: {passage #5} Question: {question} [Assistant]
+Figure 3. Prompt for RAG OpenQA where top 5 retrieved passages are used.
 4.1.2. Weakly Trained Retriever
 After the LLM weakly labels the training set, we use these labels to train DPR and ColBERT.
-For DPR, since in-batch negatives are used during training, we only need to select the top-ranked passage to curate a list of relevant question-passage pairs, thus forming 2,000 training samples.
-For all four datasets, we used two different initializaiton settings for DPR: one is using bert-base-uncased to train from scratch (denoted as “DPRbase”), and the other by resuming training from an unsupervised dense retriever using Yibin-Lei/ReContriever (Lei et al., 2023b) (denoted as “DPRReCon”).
-ColBERT’s training process differs slightly from that of DPR because it requires the sampling of negative passages.
+For DPR, since we use in-batch negatives during fine-tuning, we only need to select the top-ranked passage to curate a list of relevant question-passage pairs, thus forming 2,000 training samples.
+For all four datasets, we experiment with two different initialization settings for DPR.
+The first uses bert-base-uncased as the starting point and fine-tunes with W-RAG data from the pre-trained base BERT model (denoted as “DPRbase”).
+The second initializes from an unsupervised pre-trained dense retriever, Yibin-Lei/ReContriever (Lei et al., 2023b), and fine-tunes it with W-RAG data (denoted as “DPRReCon”).
+ColBERT’s training process differs slightly from that of DPR, as it requires sampling of negative passages.
 For each question, we select the top-ranked passage from the reranked list as the positive passage and the immediate following 10 passages as hard negatives.
-These hard negatives are relevant but not sufficiently informative to elicit the correct answer from the LLM.
+According to the definition of our weak labels, these hard negatives are relevant but not sufficiently informative to elicit the correct answer from the LLM.
 We did not explore varying the number of hard negatives, as this is not the primary focus of this work.
 Additionally, since the off-the-shelf ColBERTv2 is trained on the entire MSMARCO dataset, using it as the starting point would introduce data leakage.
-For consistency, we always train ColBERT from scratch using bert-base-uncased for all four datasets (denoted as “ColBERTbase”).
+For consistency, we always fine-tune ColBERT from the pre-trained base BERT model bert-base-uncased for all four datasets (denoted as “ColBERTbase”).
 4.1.3. OpenQA Performance
-Once the dense retriever is trained, we integrate it into the generic RAG pipeline, where it is used to retrieve the top-$K$ evidence passages for a given user question.
-The retrieved passages are directly inserted into the prompt, as shown in Figure 2 alongside the question and instructions asking the LLM to generate an answer based on the retrieved passages or, if the passages are not useful, to answer using its internal knowledge.
-For our main experiments, we restrict Llama3-8B-Instruct to generate a maximum of 20 tokens and only use top 1 retrieved passage to supplement the prompt.
+Once the dense retriever is trained, we integrate it into the generic RAG pipeline, where the retriever is used to index the evidence corpus and retrieve the top-$k$ evidence passages for a given user question.
+The top 5 retrieved passages are directly inserted into the prompt, as shown in Figure 3.
+Llama3.1-8B-Instruct is used to perform OpenQA, and we restrict the maximum output token to be 20 tokens, following (Jin et al., 2024).
 We also explored the impact of using different numbers of supplemental evidence passages on both OpenQA performance and latency, which we discuss in the ablation studies (5.4).
 We did not experiment with different LLMs for the OpenQA task, as the LLM component of RAG is not the primary focus of our study.
+Figure 4.
+Same information as Table 1, comparing the impact of retrievers fine-tuned with weak labels versus ground-truth labels on OpenQA performance, both of which are fine-tuned from the baseline model.
+ReContriever serves as the baseline for the DPR-family; ColBERT initialized from base BERT serves as the baseline for the ColBERT-family.
 4.2. Baselines
-Since our retriever is trained from scratch with limited data, we do not compare it to state-of-the-art retrievers or RAG systems.
-Instead, we focus on the improvement from untrained baselines to our weakly trained models and compare them with models trained on ground-truth data.
-For OpenQA evaluation, we use Llama3-8B-Instruct as the LLM in the RAG pipeline, assessing the impact of different retrievers on final OpenQA performance.
-We compare these retrievers against the same baseline models used for OpenQA.
-Additionally, we introduce two baselines: “Naive,” where Llama3 answers the question without any external information, showcasing its parameter knowledge, and “Groundtruth,” where the ground-truth passage is inserted into the prompt, representing the best possible OpenQA performance.
-To assess weak label quality, we compare Llama3-8B-Instruct with BM25 after Llama3 generates weak labels on the training set of each dataset.
-For retriever performance, we evaluate two categories of models.
+Since our retriever is fine-tuned with only limited data, we do not compare it to state-of-the-art retrievers or RAG systems.
+Instead, we focus on the improvement from starting point baselines to our weakly supervised models and compare them with models fine-tuned on ground-truth data.
+For OpenQA evaluation, we use Llama3.1-8B-Instruct as the LLM in the RAG pipeline, assessing the impact of the final OpenQA performance using top passages retrieved by different retrievers.
+Additionally, we introduce two baselines: “Naive,” where Llama3 answers the question without any external information, showcasing its parameter knowledge, and “Ground truth,” where the ground-truth passage is inserted into the prompt, representing the best possible OpenQA performance in our setup.
+To assess the quality of weak labels, we compare the weakly labeled ranking lists with BM25 on the training set of each dataset.
+For the fine-tuned retriever’s retrieval performance, we evaluate three categories of models.
 The first category includes entirely unsupervised models: BM25, ColBERT initialized with bert-base-uncased (denoted as “ColBERTinit”), Contriever (Izacard et al., 2021), and ReContriever (Lei et al., 2023b).
-Contriever learns by contrasting sampled passages from different documents, assuming that each document is somewhat unique, while ReContriever enhances this by estimating relevance among sampled passages.
-The second category includes DPRReCon and ColBERTbase, both trained on ground-truth data.
-For a fair comparison, we use the same questions and an equal number of positive and negative samples as training data.
-Retriever MSMARCO QnA NQ SQuAD WebQ
-F1 Rouge-L BLEU-1 F1 Rouge-L BLEU-1 F1 Rouge-L BLEU-1 F1 Rouge-L BLEU-1
-Naive 0.2749 0.2392 0.2019 0.1158 0.1136 0.0618 0.0885 0.0845 0.0508 0.1512 0.1487 0.0823
-Groundtruth 0.4677 0.4310 0.3541 0.2186 0.2171 0.1193 0.2691 0.2675 0.1613 0.1736 0.1707 0.0937
-ColBERTinit 0.2185 0.1867 0.1619 0.0525 0.0518 0.0283 0.0501 0.0482 0.0296 0.0434 0.0423 0.0233
-BM25 0.3060 0.2712 0.2292 0.1374 0.1358 0.0748 0.1387 0.1369 0.0835 0.1085 0.1055 0.0594
-Contriever 0.3125 0.2798 0.2347 0.1446 0.1432 0.0796 0.1144 0.1124 0.0680 0.1273 0.1257 0.0702
-ReContriever 0.3171 0.2854 0.2387 0.1440 0.1429 0.0789 0.1257 0.1233 0.0753 0.1274 0.1239 0.0695
-Trained on Groundtruth Data
-ColBERTbase 0.3227 0.2876 0.2421 0.1494 0.1482 0.0814 0.1524 0.1497 0.0900 0.1006 0.1001 0.0538
-DPRReCon 0.3424 0.3102 0.2588 0.1641 0.1628 0.0902 0.1550 0.1531 0.0919 0.1397 0.1378 0.0771
-Trained on W-RAG Data
-ColBERTbase 0.3150 0.2803 0.2365 0.1313 0.1301 0.0712 0.1509 0.1489 0.0897 0.1003 0.0989 0.0545
-DPRReCon 0.3397 0.3074 0.2559 0.1528 0.1515 0.0837 0.1425 0.1402 0.0850 0.1304 0.1276 0.0727
-Table 1.
-OpenQA Performance.
-Top 1 passage is inserted into the OpenQA prompt, limit number of generated tokens to 20.
-: Performance improvements are statistically significant with a ttest of $p<0.05$.
-DPRReCon’s baseline is ReContriever, ColBERTbase’s baseline is ColBERTinit
+Contriever learns by contrasting sampled passages from different documents, assuming passages sampled from the same document are more similar than those sampled from different documents.
+ReContriever enhances Contriever by adding an estimated prior of the relevance among sampled passages.
+The second category includes retrievers fine-tuned on ground-truth data, and the third category are fine-tuned with W-RAG data.
+For a fair comparison between the second and third categories, we use the same questions and an equal number of positive and negative samples as training data.
 4.3. Experimental Settings
@@
 The nltk  [^4] https://github.com/nltk/nltk tokenizer is used to tokenize the questions and passages.
-We use BEIR  [^5] https://github.com/beir-cellar/beir to manage data, and the DPR models are trained using sentence-transformers with the following hyperparameters: batch size of 128, learning rate of $2\times 10^{-5}$, AdamW optimizer, 20 epoch, and model weights are saved based on the best Recall@5 on the validation set.
-ColBERT is trained from scratch using RAGatouille  [^6] https://github.com/bclavie/RAGatouille with a batch size of 64, 10 hard negatives per positive, learning rate of $1\times 10^{-5}$, AdamW optimizer, and 1 epoch.
-Generating weak labels took approximately two days per dataset on an Nvidia V100 GPU.
-Our experiments are easily reproducible since BM25, ColBERT, DPR, Llama3, and all four datasets are publicly accessable.
-The prompts used to generate weak labels are shown in Figure 2.
-We use the term “DOCUMENT” instead of “PASSAGE” because our testing experiments showed that “DOCUMENT” consistently yields better reranking performance.
-We suspect this is because “DOCUMENT” is a more common term and likely appears more frequently in Llama3’s training data, leading it to a deeper understanding of the word “DOCUMENT”.
-Additionally, we tested different orderings of the passage, question, and instructions, finding that the Passage-Question-Instruction format consistently outperforms Passage-Instruction-Question and Instruction-Passage-Question.
+The prompt used to generate weak labels is shown in Figure 2.
+The weak relevance score between a question and a passage is extracted from the logits of the answer tokens, which is why the ground-truth answer is explicitly included in the prompt.
+We use BM25’s retrieval performance as a baseline and try different orderings of the passage, question, and instructions, finding that the weak labels generated from the “Passage-Question-Instruction” format consistently yield better retrieval metrics compared to the “Passage-Instruction-Question” and “Instruction-Passage-Question” formats.
+For fine-tuning, we borrow BEIR’s code  [^5] https://github.com/beir-cellar/beir to manage the training data.
+The DPR models are fine-tuned using sentence-transformers with the following hyperparameters: a batch size of 128, a learning rate of $2\times 10^{-5}$, the AdamW optimizer, 20 epochs, and model weights are saved based on the best Recall@5 on the validation set.
+ColBERT is fine-tuned from base BERT models using RAGatouille  [^6] https://github.com/bclavie/RAGatouille with a batch size of 64, 10 hard negatives per positive, learning rate of $1\times 10^{-5}$, AdamW optimizer, and one epoch.
+Generating weak labels took approximately two days per dataset on a single Nvidia V100 GPU.
+Our experiments are easily reproducible since BM25, ColBERT, DPR, Llama3-8B-Instruct, Llama3.1-8B-Instruct, and all four datasets are publicly accessible.
 4.4. Evaluation
-We use Recall as our primary metric to evaluate the retrieval performance, as it focuses on whether the relevant passage is included within a specific ranking range.
-We also report MRR (Mean Reciprocal Rank), which is similar to Recall but also considers the position at which the relevant passage is ranked.
-To gauge the alignment between machine-generated answers and ground-truth answers, we use F1, Rouge-L (Lin, 2004), and BLEU-1 (Papineni et al., 2002).
-These metrics collectively measure the overlap and sequence matching between generated and actual answers.
-We evaluate OpenQA performance on our uniformly random sampled test set, which contains 2,000 questions for each datasets.
-For questions with multiple correct answers in the dataset, we select the answer with the highest cumulative score across all OpenQA metrics.
-Retriever MSMARCO QnA NQ SQuAD WebQ
+To evaluate retrieval performance, we use Recall as our primary metric, as it reflects whether the relevant passage is included within a specific top-$k$ ranking range.
+We also report Mean Reciprocal Rank (MRR), which measures the inverse of the rank at which the first relevant passage appears.
+Unlike Recall, which only considers whether a relevant passage is retrieved, MRR also takes into account its position in the ranked list.
+For OpenQA metrics, we use F1, Exact Match (EM), and BLEU-1 (Papineni et al., 2002) to assess the alignment between LLM-generated answers and ground-truth answers.
+F1 measures the token-level overlap between generated and reference answers.
+EM evaluates whether the generated answer exactly matches any of the ground-truth answers.
+BLEU-1 computes unigram precision, measuring the proportion of overlapping words between the LLM-generated answer and the reference answers without considering word order or recall.
+We evaluate OpenQA performance on our randomly sampled test set, which contains 2,000 questions per dataset.
+For questions with multiple ground-truth answers, the generated answer is compared with all references, and the highest resulting score is used for evaluation.
+Retriever NQ SQuAD WebQ MSMARCO QnA
 MRR@5 R@1 R@5 MRR@5 R@1 R@5 MRR@5 R@1 R@5 MRR@5 R@1 R@5
-ColBERTinit 0.0163 0.0120 0.0220 0.0866 0.0093 0.0224 0.0504 0.0063 0.0158 0.0031 0.0001 0.0004
-BM25 0.2781 0.1647 0.4580 0.5686 0.0869 0.3575 0.2978 0.0400 0.2011 0.2737 0.0285 0.1336
-Contriever 0.2826 0.1585 0.4901 0.5524 0.0988 0.3585 0.2221 0.0294 0.1606 0.2990 0.0165 0.1304
-ReContriever 0.2908 0.1589 0.5093 0.5831 0.1118 0.3952 0.2150 0.0284 0.1597 0.3022 0.0194 0.1370
-Trained on Groundtruth Data
-$\text{DPR}_{base}$ 0.3237 0.2117 0.4994 0.5608 0.1174 0.3619 0.1786 0.0199 0.1054 0.4028 0.0561 0.2059
-ColBERTbase 0.3484 0.2097 0.5112 0.5925 0.1316 0.3082 0.2310 0.0469 0.1072 0.1157 0.0180 0.0501
-DPRReCon 0.3652 0.2284 0.5913 0.6632 0.1516 0.4715 0.2829 0.0393 0.2072 0.4155 0.0640 0.2262
-Trained on W-RAG Data
-$\text{DPR}_{base}$ 0.2981 0.1809 0.4813 0.5302 0.1012 0.3364 0.1781 0.0210 0.1020 0.3377 0.0478 0.1656
-ColBERTbase 0.3116 0.1973 0.4904 0.5229 0.0982 0.2672 0.2265 0.0445 0.1029 0.0831 0.0153 0.0326
-DPRReCon 0.3496 0.2023 0.5887 0.6301 0.1311 0.4267 0.2811 0.0382 0.2001 0.3610 0.0499 0.1880
-Table 2. Retrieval Performance after training with W-RAG or ground-truth data. : Performance improvements are statistically significant with a ttest of $p<0.05$.
-Retriever MSMARCO QnA NQ SQuAD WebQ
+Without fine-tuning
+BM25 56.86 8.69 35.75 29.78 4.00 20.11 27.37 2.85 13.36 27.81 16.47 45.80
+Contriever 55.24 9.88 35.85 22.21 2.94 16.06 29.90 1.65 13.04 28.26 15.85 49.01
+ColBERTinit 8.66 0.93 2.24 5.04 0.63 1.58 0.31 0.01 0.04 1.63 1.20 2.20
+ReContriever 58.31 11.18 39.52 21.50 2.84 15.97 30.22 1.94 13.70 29.08 15.89 50.93
+Fine-tuned with ground-truth data
+DPRbase 56.08 11.74 36.19 17.86 1.99 10.54 40.28 5.61 20.59 32.37 21.17 49.94
+ColBERTbase 59.25 13.16 30.82 23.10 4.69 10.72 11.57 1.80 5.01 34.84 20.97 51.12
+DPRReCon 66.32 15.16 47.15 28.29 3.93 20.72 41.55 6.40 22.62 36.52 22.84 59.13
+Fine-tuned with W-RAG data
+DPRbase 53.02 10.12 33.64 17.81 2.10 10.20 33.77 4.78 16.56 29.81 18.09 48.13
+ColBERTbase 52.29 9.82 26.72 22.65 4.45 10.29 8.31 1.53 3.26 31.16 19.73 49.04
+DPRReCon 63.01 13.11 42.67 28.11 3.82 20.01 36.10 4.99 18.80 34.96 20.23 58.87
+Table 2.
+Retrieval Performance after training with W-RAG or human-labeled data.
+Bold indicates the best performance; Underlined indicates the second best.
+: Performance improvements are statistically significant with a t-test of $p<0.05$.
+Retriever NQ SQuAD WebQ MSMARCO QnA
 R@1 R@5 R@50 R@1 R@5 R@50 R@1 R@5 R@50 R@1 R@5 R@50
-BM25 0.1694 0.4464 0.7577 0.088 0.353 0.7941 0.0449 0.2069 0.6461 0.0217 0.1155 0.5765
-Llama3-8B 0.5239 0.6614 0.7517 0.1542 0.4224 0.7303 0.0805 0.3088 0.6774 0.0494 0.1887 0.5502
-Table 3. BM25 retrieves the top 100 passages, which are then reranked by Llama3-8B. The reranked lists are used as weak labels for training a dense retriever.
+BM25 8.8 35.3 79.41 4.49 20.69 64.61 2.17 11.55 57.65 16.94 44.64 75.77
+Llama3-8B-Instruct 15.42 42.24 73.03 8.05 30.88 67.74 4.94 18.87 55.02 52.39 66.14 75.17
+Table 3. BM25 retrieves the top 100 passages, which are then reranked by Llama3-8B-Instruct. The reranked lists are used as weak labels for training a dense retriever.
 5. Experimental Results
-In this section we discuss the final OpenQA performance, retrieval performance of trained retrievers, the quality of weak labels, and ablation studies.
+In this section, we discuss the final OpenQA performance, the retrieval performance of fine-tuned retrievers, the quality of weak labels, and the results of ablation studies.
 5.1. Main Results
-We do not claim that W-RAG outperforms state-of-the-art RAG methods or achieves the best retrieval performance on any of the four datasets.
-Instead, we highlight the performance improvements of weakly trained retrievers compared to baselines, and demonstrate that the gap between weakly supervised and ground-truth trained retrievers is relatively small.
-This relationship is evident in both the retrieval performance and the final RAG for OpenQA performance.
-The goal of W-RAG is to enhance LLM generation quality by providing relevant evidence passages.
-Tabel 1 presents the main OpenQA results using different retrievers in the RAG pipeline, evaluated on 2,000 test questions for each dataset.
-For these experiments, only the top 1 retrieved passage is added to the prompt, and the LLM is limited to generating 20 tokens.
-The Naive and Groundtruth baselines represent the lower and upper bounds, respectively.
-The Naive baseline excludes any supplementary passage, while Groundtruth includes the best passage.
-Any retriever that retrieves relevant passages should outperform the Naive baseline, while surpassing the Groundtruth baseline would be very unexpected.
-Results show consistent improvements with W-RAG trained retrievers across all datasets, with most of them being statistically significant.
-The relatively small gap between W-RAG trained and ground-truth trained retrievers suggests that W-RAG data approaches the quality of human-labeled data.
-One baseline, ColBERTinit (initialized with bert-base-uncased), underperforms the Naive baseline, likely due to irrelevant passages being retrieved, introducing noise that misguides Llama3.
-We notice that for NQ, SQuAD, and WebQ, all the OpenQA metrics are small.
-This is due to the brevity of ground-truth answers (average length  2 words).
-For example, in response to the NQ question ”When did Big air Snowboarding become an Olympic sport?”, Llama3 outputs the correct answer: ”Big Air Snowboarding was first added to the 2018 Winter Olympics.” Since the ground-truth answer given by NQ is just ”2018”, standard OpenQA metrics like F1 and Rouge-L would yield low scores due to the denominator of the Precision term being the length of the generated answer, while the numerator is n-gram overlap which is at most 1 in this case.
-Despite this, these metrics remain valuable; if Llama3’s answer omitted ”2018,” the score would be zero.
-The absolute metric values are less important than the relative differences across settings.
-Although token length could be treated as a hyperparameter, we limited generation to 20 tokens across datasets for consistency.
+The goal of W-RAG is to draw weak training signals from the downstream task (such as OpenQA) of an LLM, and fine-tune the retriever to prioritize passages that most benefit the task.
+In other words, W-RAG aims to improve OpenQA performance by aligning retrieval with answer generation.
+We highlight the performance improvements of weakly supervised retrievers compared to baselines, and demonstrate that the gap between weakly supervised and ground-truth supervised retrievers is relatively small.
+This trend is observed in both retrieval metrics and final performance in the RAG-based OpenQA setting.
+Table 1 presents the main OpenQA results using different retrievers in the RAG pipeline, evaluated on 2,000 test questions for each dataset.
+For these experiments, the top 5 retrieved passages are added to the prompt, as shown in Figure 3.
+The “Naive” and “Ground truth” represent the lower and upper bounds of OpenQA performance under our setup, respectively.
+The “Naive” baseline excludes any supplementary passage, while “Ground truth” includes the ground-truth passage.
+Any retriever that retrieves relevant passages should outperform the “Naive” baseline, while surpassing the “Ground truth” upper bound would be very unexpected.
+Figure 4 provides a more intuitive comparison of the OpenQA results.
+Across all datasets, W-RAG-tuned retrievers consistently outperform baselines within both the DPR and ColBERT model families.
+Most of these improvements are statistically significant.
+Furthermore, the relatively small performance gap between W-RAG-tuned and ground-truth-tuned retrievers suggests that the weak labels generated by W-RAG approach the quality of human-labeled data.
+One baseline, ColBERTinit (initialized with bert-base-uncased), underperforms the “Naive” baseline, likely due to irrelevant passages being retrieved, introducing noise that misguides the LLM.
+For MSMARCO QnA, the EM values are relatively smaller because the ground-truth answers are generally much longer than NQ, SQuAD, or WebQ, and are therefore more challenging for the LLM to generate exactly the same words as the ground-truth answer.
 5.2. Retrieval Results
-The retrieval results of different trained retrievers are presented in Table 2.
-These results are based on 2,000 test questions for each trained or baseline retriever.
-On the MSMARCO QnA dataset, all retrievers trained on W-RAG data outperformed all baselines, while DPRReCon consistently outperformed all unsupervised baselines, except for SQuAD where BM25 performed best.
-As expected, DPR trained on ReContriever showed better retrieval performance than DPR trained from scratch.
-Similar to the OpenQA task, W-RAG trained retrievers slightly underperform compared to those trained on ground-truth data.
-Although there appears to be a positive correlation between retrieval performance and OpenQA performance, this relationship is noisy and sometimes inconsistent.
-For instance, W-RAG trained ColBERTbase achieved nearly 4% higher Recall@1 than ReContriever on MSMARCO QnA, ranking the relevant passage at top for about 80 more questions.
-However, when these top 1 passages are used for OpenQA, ColBERTbase performed slighly worse than ReContriever.
-Similarly, ground-truth trained DPRReCon significantly outperformed W-RAG trained DPRReCon in Recall@1 on MSMARCO QnA, yet their OpenQA performance was almost identical when using the top-ranked passage.
-These findings suggest that the traditional definition of relevance may not directly impact the answer quality in RAG systems.
-W-RAG’s top passage is selected based on its ability to elicit the correct answer from the LLM, while ground-truth passages are typically chosen based on term overlap, semantic similarity, or human judgment.
-Contrary to expectations, better retrieval does not always lead to better OpenQA.
-This raises the question: “Is the relevance definition in existing datasets truly effective for evaluating retrievers in RAG systems?”
+The retrieval results of different fine-tuned retrievers are presented in Table 2.
+These results are based on 2,000 test questions for each fine-tuned or baseline retriever.
+On NQ, WebQ, and MSMARCO QnA datasets, all retrievers fine-tuned on W-RAG data outperformed their corresponding baselines, while DPRReCon consistently outperformed all unsupervised baselines, except for SQuAD where BM25 performed best.
+As expected, DPR fine-tuned from ReContriever showed better retrieval performance than DPRbase, which was fine-tuned from the base BERT model.
+Similar to the OpenQA task, retrievers fine-tuned with W-RAG data slightly underperform those fine-tuned with ground-truth data.
+Although there is a general positive correlation between retrieval performance and OpenQA performance, the relationship is noisy and sometimes inconsistent.
+For instance, on SQuAD, BM25 achieved a much higher Recall@5 (20.11) compared to ColBERTbase fine-tuned with W-RAG data (10.29), meaning BM25 ranked the relevant passage among the top 5 for approximately 200 more questions.
+However, in OpenQA, using the top 5 passages from ColBERTbase led to better answer accuracy than using those from BM25.
+Similarly in NQ, DPRReCon fine-tuned on W-RAG achieved significantly higher retrieval performance than ColBERTbase (Recall@5 of 42.67 vs.
+26.72), their OpenQA performances were relatively close.
+This discrepancy may be partly due to the prompt instructing Llama3.1-8B-Instruct to fall back on its internal knowledge when the retrieved passage cannot answer the question.
+Nevertheless, these findings suggest that traditional notions of relevance may not directly translate into improved quality of answers in RAG systems.
+Retrievers trained on W-RAG data select top passages based on their ability to elicit the correct answer from the LLM, while ground-truth passages are typically chosen based on term overlap, semantic similarity, and human judgment.
+Contrary to expectations, stronger retrieval metrics do not always lead to better OpenQA results.
+This raises a critical question: Is the relevance definition in existing datasets truly effective for evaluating retrievers in RAG systems?
 5.3. W-RAG Labels
-After generating all of W-RAG data, we examine its retrieval performance and the quality of the weakly labeled rank lists.
-In Table 3, a significant gap between BM25 and Llama3’s reranking performance at Recall@1 is evident, which justifies our choise of only using the top 1 passage as the positive when training different retrievers.
-We also observe that Llama3’s performance declines as the number of passages increases.
-This is likely due to the fact that starting from around the 50th position, the difference in answer likelihood between adjacent passages diminishes to negligible values, such as $\exp(-3.24815)-\exp(-3.24899)\approx 3\times 10^{-5}$.
-Prompt R@1 R@5 F1 Rouge-L BLEU
-BM25 0.1343 0.4128 0.2797 0.2436 0.0640
-UPR 0.1970 0.4970 0.2898 0.2538 0.0677
-Zero-shot 0.5027 0.6289 0.4300 0.3924 0.1876
-One-shot 0.5187 0.6690 0.4358 0.3976 0.1939
-Two-shot 0.5137 0.6840 0.4256 0.3858 0.1812
-Table 4. Llama3-8B-Instruct’s reranking performance using different prompts on 500 validation questions from the MSMARCO QnA dataset. Answers are generated using only the top 1 reranked passage.
+After generating the W-RAG data, we evaluate its retrieval performance and assess the quality of the weakly labeled ranking lists.
+As shown in Table 3, there is a significant gap in Recall@1 between BM25 and the reranking performance of Llama3-8B-Instruct.
+This observation supports our decision to use only the top-1 passage as the positive example when fine-tuning different retrievers.
+We also observe that the quality of the reranked list deteriorates at higher recall values.
+This decline occurs because the difference in answer likelihood between adjacent passages becomes negligible beyond a certain rank.
+For example, the difference in answer likelihood between the 50th and 51st ranked passages is approximately $\exp(-3.24815)-\exp(-3.24899)\approx 3\times 10^{-5}$, which is very insignificant.
+For passages that do not meaningfully increase the LLM’s likelihood of generating the correct answer, the assigned likelihood scores remain similarly low across the ranking.
+Figure 5. Comparison of recall for various LLMs at different top k positions, when reranking top 100 passages retrieved by BM25.
 5.4. Ablation Study
-We assessed the quality of W-RAG data generated by different LLMs, specifically their ability to rerank passages based on answer likelihood.
-This is illustrated in Figure 3, showing the reranking performance of various LLMs on 500 questions from the validation set of MSMARCO QnA.
-Although the LLMs notably differs at Recall@5, they exhibit consistent trends with each other.
-We notice Llama3-8B-Instruct is weak when compared to other examined LLMs.
-However, since their recall distributions are similar, any of these LLMs can serve as the reranker.
-We chose to use Llama3-8B-Instruct for all our experiments.
-# F1 Rouge-L BLEU-1 Latency (sec/infer)
-1 0.3397 0.3074 0.2559 1.043
-3 0.3744 0.3395 0.2818 1.270
-5 0.3767 0.3427 0.2835 1.454
-10 0.3864 0.3524 0.2911 2.134
-Table 5. Adding different # of evidence passages to the prompt when doing OpenQA using DPRReCon on 2,000 test questions from MSMARCO QnA.
-Using Llama3-8B-Instruct, we evaluated reranking and OpenQA performance, as presented in Table 4.
-UPR (Sachan et al., 2022a), which ranks passages based on question likelihood, serve as an unsupervised approach for generating training labels.
-With weak supervision of ground-truth answers, W-RAG’s zero-shot, one-shot, and two-shot prompts significantly outperforms BM25 and UPR in both reranking and OpenQA metrics.
-However, the differences between the various shot prompts are not very consistent and are not significant.
-We hypothesize that this inconsistency may have stemmed from the LLM misinterpreting example passages as the relevant passage, or if the question types are drastically different between the example and the given question, thus introducing noise and reducing result reliability.
-Due to this, we conducted all subsequent experiments using zero-shot prompts, shown in Figure 2.
-During the final OpenQA, we also study the impact of giving more evidence passages, shown in Table 5.
-We observe a consistent and steady increase in all OpenQA metrics when more evidence passages are inserted into the prompt.
-This behavior is expected because the more passages supplied, the better chance a good passage is within them.
-However, latency also grows as we increase the number of passages.
-Figure 3. Comparison of recall for various LLMs at different top k positions, when reranking top 100 passages retrieved by BM25.
+To study the quality of W-RAG data as generated by different LLMs, we experimented with several LLMs (Jiang et al., 2023a; Rivière and et al., 2024), as shown in Figure 5.
+The figure reports reranking performance on 500 questions randomly selected from the validation set of MSMARCO QnA.
+Although the LLMs differ slightly in Recall@5, they exhibit consistent trends.
+This consistency gives us confidence that the W-RAG method generalizes well to models at the 8B parameter scale.
+We leave the more extensive study with larger models for future work.
+We notice Llama3-8B-Instruct is relatively weaker when compared to other examined LLMs.
+Nevertheless, the recall distributions across LLMs are similar, and we conclude that any of these models can effectively serve as the agent for generating weak supervision signals.
+In the 8B parameter scale, this choice does not substantially impact the outcome of the trained retrievers and the downstream OpenQA performance.
+We also examine the trade-off between OpenQA accuracy and latency on the NQ dataset by varying the number of top passages retrieved by DPRReCon, as shown in Table 4.
+We observe a consistent and steady improvement in all OpenQA metrics as more evidence passages are included in the prompt.
+This trend begins to plateau between 10 and 20 passages under our experimental setting.
+This behavior is expected: supplying more passages increases the likelihood that at least one of them contains the correct answer.
+However, latency increases linearly with the number of inserted passages, making it a trade-off between answer quality and inference speed.
+# Passages F1 EM BLEU-1 Avg Latency (ms)
+1 54.64 40.95 50.50 264.90
+3 57.74 42.90 53.14 342.01
+5 60.21 45.35 55.62 413.96
+10 61.32 46.10 56.59 602.33
+20 61.39 45.40 56.15 1005.50
+Table 4. OpenQA performance on NQ using different numbers of top passages retrieved by DPRReCon fine-tuned with W-RAG data.
 6. Conclusions and Future Work
-This paper introduces a general framework, W-RAG, for extracting weak labels from question-answer pairs to address the scarcity of training data for dense retrieval in RAG for OpenQA.
-W-RAG reranks the top-$K$ initially retrieved passages by the probability of LLMs generating the OpenQA answer conditioned on each question-passage pair.
-This ranking score computation method aligns well with the latest study (Cuconasu et al., 2024), which states that the retrieved passage should answer the question; otherwise, performance will be negatively impacted.
+This paper introduces a general framework, W-RAG, for extracting weak signals from question-answer pairs using an LLM to address the scarcity of training data for dense retrieval in RAG for OpenQA.
+W-RAG reranks the top-$k$ initially retrieved passages by the probability of LLMs generating the ground-truth answer conditioned on each question-passage pair.
+This ranking score computation method aligns well with a recent study (Cuconasu et al., 2024), which states that the retrieved passage should answer the question; otherwise, performance will be negatively impacted.
 Extensive experimental results on four public OpenQA datasets demonstrate the effectiveness of W-RAG.
-In future work, we plan to explore which types of passages most effectively enhance RAG’s performance in OpenQA, as indicated by (Cuconasu et al., 2024), where even randomly sampled tokens were beneficial in some cases.
+For future work, we plan to scale the generation of W-RAG data to a quantity comparable to that used in training state-of-the-art retrievers, and evaluate whether retrievers trained with W-RAG can surpass them in final OpenQA performance when integrated into a RAG system.
+As for future research directions, we plan to explore which types of passages most effectively enhance RAG’s performance in OpenQA, as indicated by (Cuconasu et al., 2024), where even randomly sampled tokens were beneficial in some cases.
 Understanding the types of passages preferred by LLMs will allow us to design more effective dense retrieval methods, including new structures and evaluation metrics.